### v1

# MixLLM: LLM Quantization with Global Mixed-precision between Output-features and Highly-efficient System Design

Abstract
Quantization has become one of the most effective methodologies to compress LLMs into smaller size. However, the existing quantization solutions still show limitations of either non-negligible accuracy drop or system inefficiency. In this paper, we make a comprehensive analysis of the general quantization principles on their effect to the triangle of accuracy, memory consumption and system efficiency. We propose MixLLM that explores the new optimization space of mixed-precision quantization between output feature based on the insight that different output feature matter differently in the model. MixLLM identifies the output feature with high salience in the global view rather than within each single layer, effectively assigning the larger bit-width to output features that need it most to achieve good accuracy with low memory consumption. We present the sweet spot of quantization configuration of algorithm-system co-design that leads to high accuracy and system efficiency. To address the system challenge, we design the two-step dequantization to make use of the int8 Tensor Core easily and fast data type conversion to reduce dequantization overhead significantly, and present the software pipeline to overlap the memory access, dequantization and the MatMul to the best. Extensive experiments show that with only 10% more bits, the PPL increasement can be reduced from about 0.5 in SOTA to within 0.2 for Llama 3.1 70B, while on average MMLU-Pro improves by 0.93 over the SOTA of three popular models. In addition to its superior accuracy, MixLLM also achieves state-of-the-art system efficiency.
1 Introduction
Large language models (LLMs) [7, 32] have shown remarkable performance on various tasks. But their large memory consumption and massive computation cost have become an obstacle for the efficient deployment [43, 44]. Quantization has become one of the most sufficient solution to compress LLMs into smaller size [14, 26, 45, 46], by representing the weight or activation with smaller bit-width. However, the existing quantization solutions still show limitations of either non-negligible accuracy drop or system inefficiency.
There is a triangle of characteristics for efficient LLM quantization: accuracy, memory consumption of parameters, and system efficiency of execution, which we call effectiveness triangle of quantization. The existing quantization solutions have different focus and trade-off in the triangle:
  • The weight-only methodologies target to solve the memory consumption problem, and can speedup the small-batched decoding execution that faces the memory-wall problem [43, 21]. But their accuracy drop of 4-bit quantization can be a challenge for the production workloads sensitive to accuracy, which becomes more serious in the new models with higher information density like Llama 3 [32], as illustrated in recent studies [42, 22]. Besides, the weight-only method can lead to system performance drop for large-batched workloads (e.g., the SOTA W4A16 kernel only achieves 83% performance of its float16 counterpart at batch size 512 with hidden size 4096, shown in Fig.5).
  • The weight-activation quantization represents the activation with low-bit values along with the weights, potentially lead to higher system efficiency. But it can lead to larger accuracy drop than the weight-only method as the activation is usually harder to quantize [50, 3, 27]. Besides, it introduces more dequantization overhead for the activation that can hurt the system efficiency. The transformation optimizations in some works can make the system efficiency even worse.
  • Outlier separation and mixed-precision technologies emerge to improve the accuracy of low-bit quantization by either excluding the unstructured high-salience weights from quantization [11, 21] or assigning larger bit-width for the quantization of structured high-salience weights [50]. The former shows system efficiency problem due to the low efficiency of half precision (i.e., float16/bfloat16) sparse tensor processing. The state-of-the-art mixed-precision solution [50] aims for low-bit quantization but shows non-negligible accuracy drop, even inferior to the 4-bit weight-only quantization.
Figure 1: Illustration of the quantization with mixed-precision between output features and kernel execution.
Contributions. In this paper, we provide an extensive analysis of the general quantization principles. To address the limitations of the previous works and cover the three characteristics in the effectiveness triangle, we propose MixLLM, which makes the following contributions:
[…]
Extensive evaluation shows that, with only 10% of 8-bits (i.e., W4.4A8), MixLLM outperforms all the existing 4-bit quantization algorithms while achieving the state-of-the-art system efficiency.
2 Background, Related Work, and Discussion
2.1 Background of Quantization
The quantizaiton maps the tensor $X$ into the target range with smaller bit-width representation through affine transformation: $X_{q}=clamp(\lfloor\frac{X}{s}\rceil+z,range)$, where $s$ is the scale and $z$ is the zero point. The value can be recovered (i.e., dequantization) through: $X^{\prime}=(X_{q}-z)\times s$. $X^{\prime}$ is pushed to the discrete chunks rather than recovered to the original value, thus has accuracy loss. The bit-width is essential for the accuracy of quantization as it determines the number of chunks for the quantized values ($2^{bit\_width}$). Take an example, enlarging the bit-width from 4 to 5 can double the number of chunks, so that the 5-bit RTN quantization can easily beat the 4-bit quantizations of advanced techniques (Tab.1).
The scale and zero point can be calculated from the whole channel/token vector or a small group within the channel/token, the former is called per-channel/token quantization and the latter is group-wise quantization. The group-wise scheme results in smaller accuracy loss due to the smaller chunk scale, but requires more complex GPU kernel design. The symmetric quantization uses 0 as the zero point value, which simplifies the computations ($X_{q}=clamp(\lfloor\frac{W}{s}\rceil,range)$, $X^{\prime}=X_{q}\times s$) and enables many works to design the per-channel/per-token quantized kernels by multiplying the scales at the epilogue of the whole MatMul (matrix multiplication) for dequantization [45, 40]. However, it leads to larger loss than the asymmetric one as the data distribution can be usually asymmetric, especially for smaller bit-widths like 4-bit.
2.2 Related Works and Discussion of General Quantization Principles
This paper mainly focuses on post-training quantization (PTQ).
Systems that affect the quantization requirement. The continuous batching technology [47] enables to batch the decoding tasks from different requests together to enlarge the batch dimension of MatMul during LLM inference. The chunked-prefill method [19, 2, 51] advances the continuous batching by merging the prefill and decoding tasks into the same batch, further enlarging the MatMul shapes. These technologies pushes many LLM jobs to become compute-bound and motivate the demand to reduce computation.
Weight-only quantization and its limitation. There emerges a wide range of technologies to improve the accuracy of weight-only quantization. GPTQ [14] advances OBC [13] on OBS-based [18] weight compensation with blocked updating and reordering. AWQ [26] proposes to scale the weight according to the characteristic of activation. OminiQuant [37]) proposes the learnable scaling and weight clipping factors. SpQR [11], SqueezeLLM [21] and OWQ [25] separate the outliers from the quantiation and with half precision. QuiP [8] aims to achieve extreme low-bit quantization with incoherence processing. ZeroQuant(4+2) [42] aims to improve accuracy with medium-sized FP6 quantization.
The weight-only quantization does not reduce the computation but introduces the extra dequantization operations. The low-bit weight will be dequantized to float16 to execute the MatMul in float16 datatype. The current weight-only quantization faces two challenges: 1) From the accuracy aspect, there is still an accuracy gap between the 4-bit quantization and the float16 model, especially for many real business scenarios sensitive to the small accuracy drop, as discussed in the recent works [42, 44]. 2) It can lead to system efficiency drop on busy servers as the recent LLM inference serving systems will usually batch the processing of different requests together on the server and form large MatMuls. The large MatMuls are compute-bound and will suffer from the dequantization overhead [27].
Weight-activation quantization and the challenges. The weight-activation quantization helps to make use of the low-bit computing unit. LLM.int8() [10] observes the activation outlier problem and separates outliers from quantization with half precision. ZeroQuant [46] proposes the per-token activation quantization and group-wise weight quantization. SmoothQuant [45] addresses the activation outlier problem through smoothing, and AffineQuant [29] proposes the general affine transformation for quantization. RPTQ [48] reorders the channels to cluster similar scaled values together. SpinQuant [28] and QuaRot [3] leverages matrix rotation properties to alleviate the outlier phenomenon. Atom [50] uses the mixed-precision between input features to improve accuracy of 4-bit activation quantization. QoQ [27] is a holistic quantization solution with progressive group quantization, attention smoothing, rotation, and channel reordering.
Even though the weight-activation quantization has the advantage of reduced MatMul computation (i.e., MatMul in smaller bit-width to make use of the smaller bit-width computing unit with higher computation throughput [^2] The extra dynamic activation quantization kernel can be fused into other operators with very little system cost [50], thus we only discuss the MatMul itself.), it faces the challenge of accuracy drop caused by the activation quantization, especially that the activation is usually harder to quantize than the weight. The SOTA low-bit weight-activation solutions [3, 28, 27] still have a gap to the 4-bit weight only quantization.
Beside the accuracy drop, the activation quantization will introduce more dequantization overhead than the weight-only one, which makes it challenging to design efficient GPU kernels. When enabling the asymmetric quantization, the result of $(X_{q}-z)$ may exceed the range of the bit-width of $X_{q}$, making it hard to use the corresponding Tensor Core computation. Systems like Atom [50] thus avoid using the asymmetric quantization, with the cost of larger accuracy drop. The group-wise quantization requires fine-grained integer-to-float (I2F) conversion to apply per-group scales. However, the I2F instruction is more expensive than the common computation instructions on the GPU [1] and can lead to large system performance drop ($>10\%$ drop in our practice). Besides, the throughput of Tensor Core is much higher than that of SIMT Cores, 624 TOPS of int8 Tensor Core vs. 19.5 TFLOPS/TOPS of FP32/INT32 SIMT Cores. There still lacks a well designed software pipeline to overlap the Tensor Core computation and SIMT Core based dequantization in the existing works while achieving a high accuracy.
In general, the existing solutions focus on partial of the effectiveness triangle, but cannot cover all of them well. MixLLM is orthogonal to the above works by exploring the mixed-precision between output features with global salience identification, and the co-designed quantization decision and GPU kernels.
3 Methodology
3.1 Quantization Design and Decision in MixLLM
To cover the three aspects of the effectiveness triangle simultaneously, we make the following design and decision of weight and activation quantization according to the analysis in Sec.2.2.
3.1.1 Mixed-precision between output features of weight, with global salience identification.
It is known that different elements of the weight show different salience to the network’s loss when being quantized [21, 11]. The outlier separation method can improve the accuracy by using float16 to store the high-salience elements, but can suffer from the inefficient sparse MatMul. We observe that the elements with high salience tend to show distribution along the output channels for most of the linear layers in many LLMs. Based on this observation, we can assign larger bit-width to the output channels of high salience, and smaller bit-with to the others, forming structured mixed-precision quantization. Through the experiments, we get the same conclusion with the existing works [21, 11] that there is only a small set of elements with high salience contributing significantly to the model’s accuracy drop. Thus we only need to assign the large bit-width to a small portion of the output channels to achieve good accuracy and retain a small memory consumption at the same time.
The structured mixed-precision between different output channels can be friendly to the system efficiency and kernel development, due to the nature that different output features are disjoint in the MatMul and the computation of them are different sub-problems. Fig.1 shows how the linear layer computes with the mixed-precision between output features. It divides the linear into independent sub-problems, and finally gathers the output of the sub-problems together to form the final result. This optimization space is orthogonal to the existing quantization optimizations, e.g., GPTQ [14], and can be applied together with them.
One critical problem is how to identify the high-salience output channels in the model. The fixed threshold [11] or the fixed number/ratio [50, 25] of high salience elements computed by the local loss of layers can be sub-optimal to the end-to-end model, as different layers can show different importance to the model’s final output [16, 30, 12]. A high salience channel w.r.t. a layer may not be a high salience channel of the end-to-end model. In MixLLM, we compute the high salience channels globally according to their impact to the model’s final loss (Sec.3.2). As a result, different layers will have different number of high salience channels. Fig.2 shows the distribution of the top 10% high-salient out features in Llama 3.1 8B, showing huge difference in different linear layers.
Figure 2: The percentage of high-salient out features within each linear layer of Llama 3.1 8B model according to each feature’s contribution to the final loss after quantizing to 4-bit, with 10% high-salient features globally. Each decoder layer contains q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj in order.
Note that this design is different from the mixed-precision in Atom [50] from two aspects. 1) MixLLM first addresses the problem of identifying the high-salience channels globally rather than locally. 2) MixLLM applies the mixed-precision between output features rather than input features, which is more system performant and algorithm flexible [^3] One example is that, the outlier number in each MatMul should be a multiplier of the corresponding tiling size of kernel design to achieve a good system efficiency in Atom, which limits the flexibility of algorithm. as the output features are disjoint naturally.
3.1.2 Quantization decision with algorithm-system consideration: 8-bit symmetric activation and 4-bit asymmetric weight quantization in group-wise manner.
MixLLM makes the same decision with QoQ [27] on activation quantization to use 8-bit, as the 4-bit activation can lead to a large accuracy drop but does not lead to significant system efficiency improvement as MatMul execution tends to be bound more on the larger weight tensor rather than the smaller activation tensor. It can be partially indicated from the compute intensity of the linear layer. Given token number $M$ and in/out features $K$/$N$, the compute intensity $I=\frac{2MNK}{MKB_{act}+KNB{weight}}$. $B_{act}$ and $B_{weight}$ are the bytes per element of activation and weight. Given $M=512$ and $N=K=4096$, reducing $B_{weight}$ from 8 to 4 will results in an 80% increasement of $I$, while reducing $B_{act}$ from 8 to 4 will only achieve 5.88% increasement.
Instead of using per-token activation quantization, MixLLM uses group-wise RTN method. On the one hand, Tab.1 shows that simple group-wise RTN quantization performs better than token-wise smoothing method. On the other hand, the weight is already group-wise in MixLLM, and the group-wise activation does not lead to significant more dequantization overhead in the system. We observe symmetric quantization is enough for the 8-bit activation (refer to MixLLM W8A8 in Tab.1), while asymmetric can be essential for the 4-bit weight. The group-wise method with asymmetric can lead to a difficulty for the kernel to make use int8 Tensor Core, for which QoQ [27] introduces the two-step quantization method. Instead, we design a two-step dequantization method with the property of the mix of symmetric and asymmetric (Sec.3.3).
3.2 Global Precision Search Algorithm
As discussed in Sec.3.1, MixLLM determines the precision of all output features in all layers globally rather than locally. It identifies the salience of these features with respect to the final loss of the model, and assigns larger bit-width to the features leading to larger loss.
Specifically, it calculates the salience $S$ of a channel $c$ as:
$S_{c}=\lvert l(c_{q})-l(c_{0})\rvert$ (1)
which is the distance of the model’s loss between quantizing and not quantizing this single channel. In Eq.1, $l()$ is the loss function of the model w.r.t. a single channel, $c_{q}$ is the quantized weight of the channel and $c_{0}$ is the original weight. Note that it regards other neurons except $c$ as constant in $l()$.
We use the Taylor Expansion method to estimate the loss function $l(c)$ (similar with the existing quantization works, ignoring the high-order items):
$l(c)\approx l(c_{0})+g^{T}(c-c_{0})+\frac{1}{2}(c-c_{0})^{T}H(c-c_{0})$ (2)
where $g=\mathbb{E}[\frac{\partial}{\partial c}l(c)]$ is the loss’s gradient w.r.t. the channel, and $H=\mathbb{E}[\frac{\partial^{2}}{\partial c^{2}}l(c)]$ is the second-order gradient (i.e., Hessian matrix) w.r.t. the channel.
It is infeasible to calculate the Hessian matrix as it is too costly. We approximate the Hessian $H$ with the (empirical) Fisher information matrix $\mathit{F}$ on the calibration dataset $\mathit{D}$:
$H\approx\mathit{F}=\frac{1}{\lvert\mathit{D}\rvert}\sum_{d\in\mathit{D}}{g_{d}g_{d}^{T}}$ (3)
Note that $\mathit{F}$ is w.r.t. a channel, differing from the diagonal Fisher information matrix in the recent works that ignores any cross-neuron interactions [23, 21].
Based on this approximation, the second order loss factor $\frac{1}{2}(c-c_{0})^{T}(g_{d}g_{d}^{T})(c-c_{0})$ can be further simplified to $\frac{1}{2}(g_{d}^{T}(c-c_{0}))^{2}$, simplifying the expensive chained matrix multiplication into a single vector product. Finally, the salience can be calculated by:
$S_{c}=\frac{1}{\lvert\mathit{D}\rvert}\sum_{d\in\mathit{D}}\lvert g_{d}^{T}(c_{q}-c_{0})+\frac{1}{2}(g_{d}^{T}(c_{q}-c_{0}))^{2}\rvert$ (4)
We do not ignore the first order information during the calculation, differing from OBD [24] and many recent quantization works [14, 11, 21]. This is because the first order factor can be more significant in the estimation in Eq.4, as the estimated second order factor is the square of the first order factor divided by two for each sample. Considering that $g$ can be very small for the well pretrained models and the delta of the quantized weight is usually not large, the first order factor can be larger than the second order one. Besides, what we require is the loss itself rather than the arguments of the loss function, and thus we do not need to ignore the first order factor to simplify the arguments calculation.
Algorithm 1 Global precision search procedure.
1:Linear layer number $L$, weight and gradient of all linear layers ($W_{i}\in\mathbb{R}^{O,I}$, $G_{i}\in\mathbb{R}^{O,I}$ for layer $i$).
2:Global channel index with large and small bit width (largebit_channels, smallbit_channels).
3:$S_{global}$leftarrow []
4:for $i=1,2,...,L$ do
5:     $W_{i}^{delta}$leftarrow quantize($W_{i}$) - $W_{i}$
6:     $S_{1st}$leftarrow sum($G_{i}\odot W_{i}^{delta}$, dim=1) $\triangleright$ Per-channel dot product between $G_{i}$ and $W_{i}^{delta}$
7:     $S_{2nd}\leftarrow 0.5*(S_{1st})^{2}$
8:     $S\leftarrow\lvert S_{1st}+S_{2nd}\rvert$triangleright $S\in\mathbb{R}^{O}$, the salience of the $O$ channels
9:     for $channel\_id=1,2,..,O$ do $\triangleright$ Log the salience of each output channel of this layer
10:         $S_{global}$.append(tuple($i$, $channel\_id$, $S[channel\_id]$))
11:sort($S_{global}$) $\triangleright$ Sort according to the salience, descending
12:largebit_channels, smallbit_channels $\leftarrow$ $S_{global}[:N_{largebit}]$, $S_{global}[N_{largebit}:]$
Algo.1 illustrates the procedure of the global precision search. It calculates the salience of all the output channels of all linear layers and sort them in descending order globally. Given the global threshold $N_{largebit}$ as the number of large-bit precision channels, the first $N_{largebit}$ channels are intended to be quantized with 8-bit, and the other channels will be quantized with smaller bit-width (i.e., 4-bit in this paper). Any quantization methodologies (e.g., GPTQ, clip search) can be applied independently to these two disjoint parts of channels. Note that we calculate the salience of the channels in one pass rather than iterative identifying the high-salience parts in a smaller step, as we observe the single-pass method show similar results with the iterative method and saves a lot of computation overhead than the latter.
3.3 Efficient Quantization Computation System
Two-step dequantization to make use of int8 Tensor Core. As for the W4A8 computation, the dequantized weight and activations are $(W_{q}-z)s_{w}$ and $A_{q}s_{a}$, where $W_{q}$ and $z$ are uint4 datatype, $A_{q}$ is int8 datatype, and $s_{w}$ and $s_{a}$ are float16 datatype. Directly dequantizing the tensors into float16 datatype before the MatMul computation will prevent us using the fast 8-bit Tensor Core on the GPU. Instead, MixLLM uses a two step dequantization within each group. Specifically, MixLLM first partially dequantizes the weight into $(W_{q}-z)$, and then multiply it by $A_{q}$ with the 8-bit Tensor Core. Finally, it multiplies this MatMul result by the two scales within each group. Note that we use int8 datatype for $(W_{q}-z)$ so that there is no overflow problem.
Figure 3: The float and integer value of binary (010010110xx...x), each within a consecutive range.
Fast Int to Float conversion with partially fusing into Tensor Core instruction. In the above two-step dequantization computation, the step 2 is the MatMul between the integer tensor $A_{q}(W_{q}-z)$ and the float tensor $s_{a}s_{w}$. It requires the integer to float conversion (I2F) before the multiply operation. The I2F instruction is expensive on the modern GPUs. Instead, we make use of the range-dependent fast I2F transformation to convert the I2F instruction into two add/sub instructions [^4] CUTLASS [9] also has an implementation of fast I2F for general purpose.. Specifically, it is based on the fact that there exists a certain range where an integer value’s binary is the same as a corresponding float binary. As shown in Fig.3, the binary with the first 9 bits as 010010110 represents a series of consecutive int32 and float32 values, respectively. We can add a bias to an integer value to make it within this consecutive range, and then subtract a corresponding bias in float (same underling binary) to restore its value in float type. We take the middle value in this range as the bias to maximize the data range that can be safely converted, whose hexadecimal number is 0x4b400000 (i.e., in the remaining 23 bits, the first bit is 1 and the other bits are 0). This allows to convert a consecutive range of $2^{23}$ int32 numbers to float32. The range of dot product of $k$ int8 elements is $2^{16}k$, thus the above fast I2F conversion allows the $k$ value up to 128. We use quantization group size as 128 and can use the fast I2F safely:
1// bias_int = as_int(0x4b400000), bias_fp = as_float(0x4b400000);
2int tmp = src_int + bias_int;
3int dst_float = *((float*)&tmp) - bias_fp;
We further fuse the integer subtraction into the Tensor Core mma (Matrix Multiply-Accumulate) instruction. The mma instruction computes $D=AB+D$ during the MatMul computation. We initialize the accumulator $D$ as the bias_int before MatMul computation of each quantization group, and will only need to subtract the bias_fp after the MatMul. In another word, the expensive I2F is converted into a single float subtraction. The above I2F simplification brings more than 20 TOPS performance improvement for the 512/4096/4096 (M/N/K) quantized MatMul computation on an A100 GPU.
Figure 4: The GPU kernel software pipeline of group-wise W4A8/W8A8 quantized MatMul. It assumes perfect overlapping. G2S: load global to shared memory; S2R: load shared memory to register; MMA: matrix multiply-accumulation; I2F: integer to float conversion; deq: dequantize; acc: accumulate.
End-to-end software pipeline of the quantized linear kernel on the GPU. Fig.4 shows the software pipeline of the quantized kernel. Beside the basic warp tile and block tile, we introduce the quantization group tile for the fast I2F and per-group scale multiplication. It uses two output buffers for the output accumulation at register level, one for the per-group accumulation, and the other for the global accumulation. This allows to apply the per-group scales on the group-level buffer. We initialize the group buffer with the bias_int at the beginning of the group tile, and subtract bias_fp at the end of the group tile. As for the two-step dequantization, the first step is within the warp tile where each input element will subtract the zero-point before feeding into MMA, the second step is at the end of the group tile by multiplying the scale. We use the vectorized intrinsic to perform four int8 subtract in a single instruction (vsub4[27]. Besides, to improve the global memory loading efficiency, we prepack the memory layout of the weight tensor ahead-of-time to avoid the runtime permutation of the input elements. In general, this software pipeline can overlap the memory loading and computation, and the dequantization computation with SIMT Core and the MatMul computation with Tensor Core to the best, and minimizes the overhead of group-wise dequantization.
Parallel execution of sub-problems of different bit-width. As for the execution shown in Fig.1, MixLLM executes different sub-problems in parallel on the GPU with CUDA Graph. Finally, the MatMul execution of the two parts write to the same target tensor with different channel indices to generate the final output. We implement this function with the fused epilogue of MatMul to scatter the output to the corresponding indices, which is basically costless.
4 Evaluation
4.1 Setup
As for MixLLM evaluation in this paper, we use 0%, 10%, 20%, 50% and 100% percent of 8-bit based on the 4-bit quantization, respectively. Meanwhile, we use 8-bit for activation quantization. Both the weight and activation are group-wise quantized with group size 128. The 4-bit part is asymmetric quantized and the 8-bit part (including that in weight) is symmetric, which is a good trade-off between accuracy and system efficiency. Note that any other bit-width percentage configuration can be used for real scenarios to trade-off memory usage, system efficiency and accuracy in practice. We enable GPTQ in MixLLM for the all models. We also apply clip search for all the models except for the 32B, 70B and 72B models as the clip search is time consuming for large models.
Baselines and configurations. We compare MixLLM with the state-of-the-art (SOTA) quantization solutions of both weight-only and weight-activation methods. As for the weight only quantization, we compare MixLLM with the basic round-to-nearst (RTN) 4-bit and 5-bit quantization, and the production-level SOTA GPTQ [14] and AWQ[26]. As for the weight-activation quantization, we compare MixLLM with the most widely used SmoothQuant [45] and the recent SOTA QoQ [27] and QuaRot[3] (of both W4A4 and W4A8). The 8-bit tensors are all symmetric quantized in all baselines and MixLLM. We also compare the perplexity with SqueezeLLM[21], OminiQuant[37], AffineQuant[29], Atom[50] and SpinQuant[28] according to their reported numbers.
We make use of AutoGPTQ lib [5] (v0.7.1) to evaluate GPTQ, AutoAWQ lib [4] (v0.2.7) to evaluate AWQ, lmquant lib [33] (commit 9d62c5c) to evaluate SmoothQuant and QoQ, and the official repo to evaluat QuaRot. We enable the reorder trick for GPTQ evaluation, and use asymmetric and group size 128 for both GPTQ and AWQ. We follow the official configurations in lmquant to use 0.85/0.15 as the alpha/beta parameter for SmoothQuant, and 0.3/0.7 for QoQ. We use symmetric and per-channel/token quantization in QuaRot, following the configuration in its paper. We disable the KV quantization of QoQ and QuaRot in our experiments to make the comparison fair.
Models and Datasets. We evaluate MixLLM and the baselines on a variety of widely used LLMs of different sizes, ranging from 0.5B to 72B. The models include Llama 3.1 8B and 70B [32], Llama 3.2 1B, Qwen2.5 0.5B, 1.5B, 7B and 32B [17], and Mistral 7B v0.3 [20].
We use wikitext2 dataset [31] as the calibration set for GPTQ and MixLLM. We use the default pile dataset [34] as the calibration dataset for AWQ, SmoothQuant and QoQ, to enable their better performance. GPTQ, AWQ and MixLLM uses 128 samples with sequence length of 2048 for calibration. SmoothQuant and QoQ uses 64 samples with sequence length of 1024 for calibration (larger dataset results in OOM in our experiment).
Metrics. As for the algorithm accuracy, we compare the perplexity (ppl) between all the baselines on wikitext2 and C4 [35] dataset. Meanwhile, we compare a set of popular downstream tasks on Llama 3.1 8B, Qwen2.5 7B, and Mixtral 7B v0.3, including BBH [39], GPQA [36], MMLU-Pro [41], MuSR [38], ARC challenge [6], and HellaSwag [49]. We use lm-eval [15] for the downstream tasks evaluation, for which the task names are leaderboard_bbh, leaderboard_gpqa, leaderboard_mmlu_pro, leaderboard_musr, arc_challenge, and hellaswag. We use the average number for each of the tasks.
We conduct the system experiments on NVIDIA A100 (80G) GPUs with CUDA 12.1. We use PyTorch 2.4.1 and transformers 4.45.2.
4.2 Perplexity Evaluation
Table 1: Perplexity evaluation ($\downarrow$) on wikitext2/c4 (gray for c4), sequence length 2048. NA means no support. Abn means the value is too large ($>10^{5}$). For MixLLM, pn means $n\%$ 8-bit.
Llama 3.1/3.2 Qwen2.5 Mistral
baselines 1B 8B 70B 0.5B 1.5B 7B 32B 7B v0.3
float16 9.75/12.72 6.24/8.95 2.81/6.68 13.07/17.55 9.26/13.11 6.85/10.44 5.02/8.95 5.32/7.84
W4A16 11.72/15.56 6.82/9.72 3.55/7.43 15.54/20.55 10.35/14.35 7.23/10.88 5.27/9.14 5.51/8.04
RTN W5A16 10.15/13.25 6.40/9.15 3.16/9.52 13.61/18.17 9.52/13.38 6.95/10.53 5.09/8.99 5.38/7.91
GPTQ W4A16 10.38/14.15 6.52/9.55 Abn/Abn 14.01/19.04 9.64/13.75 7.09/10.75 5.20/9.08 5.49/8.19
AWQ W4A16 10.81/14.12 6.65/9.48 3.28/6.96 15.04/19.75 9.95/13.85 7.10/10.71 5.23/9.08 5.44/7.98
SmoothQuant W8A8 9.89/12.91 6.34/9.08 2.92/6.77 13.84/18.40 9.63/13.49 7.17/10.85 5.12/9.04 5.35/7.88
QoQ W4A8 Abn/Abn 6.64/9.49 3.49/7.07 Abn/Abn Abn/Abn 7.39/11.06 5.55/9.31 5.44/7.98
W4A4 Abn/Abn 8.34/11.95 6.16/9.91 NA/NA Abn/Abn 8.15/12.05 6.26/9.98 5.83/8.50
QuaRot W4A8 Abn/Abn 6.60/9.67 3.43/7.10 NA/NA Abn/Abn 7.03/10.68 5.23/9.10 5.40/7.99
W4A8 (p0) 10.36/14.09 6.54/9.62 3.30/7.24 14.43/19.61 9.66/13.79 7.03/10.75 5.21/9.08 5.42/8.02
W4.4A8 (p10) 10.05/13.51 6.42/9.33 3.02/6.83 13.42/18.13 9.44/13.43 6.92/10.57 5.12/9.01 5.36/7.93
W4.8A8 (p20) 9.95/13.25 6.37/9.22 2.97/6.79 13.32/17.99 9.40/13.35 6.90/10.53 5.09/9.00 5.35/7.90
W6A8 (p50) 9.85/12.98 6.30/9.09 2.86/6.73 13.21/17.78 9.33/13.25 6.88/10.49 5.05/8.98 5.33/7.87
MixLLM W8A8 (p100) 9.76/12.75 6.25/8.97 2.81/6.68 13.12/17.60 9.28/13.14 6.86/10.45 5.02/8.96 5.32/7.84
Tab.1 shows the perplexity on Wikitext2 and C4 dataset for the commonly used open source LLMs, of different baselines. It shows that:
  • Using 4.4 bits of weights with MixLLM can achieve the similar accuracy with 5 bits RTN weight-only quantization, even with 8-bit activation quantization enabled in MixLLM. This is mainly because MixLLM assigns the high-salience output channels with larger bit-widths than the uniform 5-bit solution.
  • As for the weight-only quantization baselines, MixLLM W4.4A8 outperforms the production SOTA solutions GPTQ and AWQ, with only 10% more bit-width, and even with 8-bit activation quantization enabled in MixLLM. Meanwhile, the RTN W5A16 method also outperforms GPTQ and AWQ, which means a slightly larger bit-width can defeat the well tuned smaller bit-width easily. MixLLM W4.4A8 benefits from the larger bits on the top 10% output features with high salience.
  • As for the weight-activation quantization baselines, MixLLM W4.4A8 shows a comparable accuracy with SmoothQuant with much smaller bit-width (60% of that in SmoothQuant). MixLLM W4.4A8 shows better accuracy than QoQ and QuaRot with only 10% larger bit-width. It shows MixLLM achieves a good balance of memory consumption and accuracy.
  • Note that MixLLM W8A8 quantization shows nearly lossless performance compared to the float16 baseline. This is part of the motivation that MixLLM uses group-wise quantization for the activation.
Table 2: PPL (wikitext2) comparison with the reported numbers in the related works.
LLaMA 2 FP16
SqueezeLLM
W4A16 0.45%
OminiQuant
W4A16/W4A4
AfineQuant
W4A16/W4A4
Atom
W4A4 128 outliers
SpinQuant
W4A8
MixLLM
W4.4A8
7B 5.47 5.57 5.58/14.26 5.58/12.69 6.03 5.7 5.55
13B 4.88 4.96 4.95/12.30 4.95/11.45 5.27 5.0 4.93
Comparison with More Related Works We compare MixLLM with more recent quantization works according to the reported numbers in their papers (Tab.2), showing that MixLLM achieves superior accuracy to a broad range of related works with similar memory consumption.
4.3 System Performance
We have evaluated MixLLM for the single linear layer of token number ranging from 1 to 1024 with in features 4096 and out features 4096/14336, and compared it with the SOTA W4A16 (TRT-LLM) and QoQ [27], shown in Fig.5. It also shows MixLLM kernel performance of different percent of 8-bits (W4A8 0% 8-bit, W4.4A8 10% 8-bit, and W8A8 100% 8-bit). It shows that:
  • MixLLM outperforms the float16 counterpart for all token numbers, achieving 1.90$\times$, 2.75$\times$, and 1.88$\times$ averaged speedup with MixLLM W4A8, W8A8, and W4.8A8 respectively.
  • MixLLM outperforms the SOTA W4A16 solution, achieving 1.26$\times$, 1.78$\times$, and 1.25$\times$ averaged speedup with MixLLM W4A8, W8A8, and W4.8A8 respectively.
  • MixLLM achieves similar performance with QoQ with similar bit-width, achieving 0.99$\times$, 1.39$\times$, and 0.99$\times$ averaged speedup with MixLLM W4A8, W8A8, and W4.8A8 respectively. Note that MixLLM has better accuracy than QoQ (Tab.1, Tab.3).
Figure 5: The speedup of two types of single linear layers over torch float16 baseline on the A100 GPU.
4.4 Downstream Tasks Evaluation
Table 3: Downstream tasks evaluation ($\uparrow$) on Llama-3.1-8B/Qwen2.5-7B/Mistral-7B-v0.3. The above is the average of the three models. BBH is 3 shot, MMLU pro is 5 shot, and others are zero shot.
BBH GPQA MMLU-Pro MuSR ARCc HellaSwag
float16
48.62
46.52/54.09/45.25
30.86
31.08/33.11/28.39
35.52
32.91/43.86/29.80
41.07
37.99/44.51/40.72
52.24
53.41/51.02/52.30
79.43
78.92/78.94/80.43
GPTQ W4A16
47.08
46.28/51.97/43.00
30.98
30.81/34.59/27.53
33.66
31.20/41.02/28.75
41.83
39.06/44.64/41.78
51.37
53.24/48.98/51.88
78.46
77.85/78.10/79.44
AWQ W4A16
47.62
45.59/52.71/44.57
30.86
29.76/33.17/29.65
34.31
31.57/42.78/28.59
40.36
38.27/42.37/40.44
51.17
52.39/50.60/50.51
78.77
78.29/78.35/79.68
SmoothQuant W8A8
47.82
46.37/52.57/44.52
30.90
31.40/33.94/27.36
35.04
32.61/42.98/29.52
42.06
39.05/46.39/40.73
51.74
53.33/50.00/51.88
79.20
78.88/78.48/80.24
QoQ W4A8
45.78
40.98/51.23/45.14
30.02
28.99/32.50/28.56
32.84
28.16/41.72/28.63
39.92
37.60/41.59/40.57
50.54
51.28/49.15/51.19
78.10
76.90/77.52/79.89
Quarot W4A4
41.10
36.96/45.42/40.92
27.53
25.41/28.94/28.23
27.60
22.99/34.40/25.42
39.46
37.92/40.68/39.77
45.99
43.00/46.33/48.63
74.85
72.87/73.54/78.14
Quarot W4A8
46.95
44.95/52.98/42.92
30.28
30.96/30.71/29.18
33.60
29.95/42.45/28.41
41.65
39.05/45.58/40.32
51.39
50.00/52.30/51.88
78.55
77.83/77.84/79.98
MixLLM W4A8
46.92
43.44/44.75/52.59
29.90
29.58/28.26/31.87
33.75
30.18/29.59/41.49
41.70
38.81/43.11/43.19
51.82
51.71/51.88/51.88
78.61
77.94/79.71/78.17
MixLLM W4.4A8
48.17
46.27/52.58/45.66
30.09
29.17/31.75/29.36
34.53
31.08/43.26/29.26
41.74
39.32/44.79/41.11
52.70
53.67/51.96/52.47
79.00
78.20/78.58/80.21
MixLLM W8A8
48.84
46.84/54.35/45.34
30.93
30.51/33.21/29.07
35.54
33.00/43.80/29.83
40.94
37.32/44.91/40.59
52.10
53.24/50.94/52.13
79.42
78.98/78.88/80.40
Tab.3 shows the accuracy of the downstream tasks on three popular LLMs. The result shows that:
  • MixLLM W4.4A8 outperforms all the 4-bit weight quantizations, with only 10% more weight consumption. For example, for the MMLU-Pro task, the average metric of MixLLM W4.4A8 is improved by 1.69, 6.93, and 0.93 over QoQ, QuaRot W4A4, and QuaRot W4A8, respectively.
  • MixLLM W8A8 is nearly lossless, showing higher accuracy than SmoothQuant. This comes from the group-wise quantized activation of MixLLM.
4.5 Detailed Analysis
4.5.1 Ablation Study
Figure 6: The perplexity (wikitext2) of Llama 3.1 8B model with different configurations.
Fig.6 shows the perplexity of Llama 3.1 8B model by adding different optimizations gradually. With the basic RTN quantization, using 8-bit for activation, and asymmetric and group-wised weight quantization contribute significantly to the accuracy improvement. This demonstrates the effectiveness of the decisions made in Sec.3.1.2. Based on these decisions, the 10% of 8-bit output features improves the accuracy to a high level, for which using blocked Fisher and including the first-order Taylor factor also contributes to the accuracy. Finally, applying GPTQ and clipping can further improve the accuracy.
4.5.2 High Precision Distribution
Fig.2 shows the percentage of 8-bit out features in each of the linear layers of Llama 3.1 8B, with 10% global 8-bit out features searched by MixLLM (i.e., W4.4A8). It shows that high-salient (i.e., 8-bit) features are distributed very differently in different linear layers. Specifically, the v_proj and down_proj layers show much higher percentage of high-salient features than other layers, for which Tab.4 shows the average percentage of different classes of linear layers.
Table 4: The average percentage of 8-bit out features in the seven classes of linear layers in Llama 3.1 8B, with 10% global 8-bit out features in MixLLM.
layer q_proj k_proj v_proj o_proj gate_proj up_proj down_proj
avg 8-bit percent (%) 3.93 12.36 71.22 18.70 0.73 1.46 53.82
4.5.3 One-pass vs. Progressive Search
As described in Sec.3.2, MixLLM searches the high-salience features within a single pass. We have tried the progressive procedure on Llama 3.1 8B and Mistral 7B models, which identifies smaller portions of the high-salience features iteratively. Results show that the accuracy is the same to the one-pass method to two decimal places. However, the progressive method shows much higher search time due to the repeated procedure. The one-pass method takes 7 minutes for each of the two models to search 10% high-salience features, while the progressive method that searches 2% high-salience iteratively takes 30 minutes to find top 10% features.
4.5.4 Overhead of Global Precision Search
Table 5: The overhead of global precision search in MixLLM.
Models Llama 3.1 Mistral Qwen2.5
8B 70B 7B v0.3 1.5B 7B
Time (min) 7 55 7 2 7
Tab.5 shows the global precision search overhead described in Sec.3.2. As noted in Sec.4.1, the calibration dataset has 128 samples with sequence length of 2048. We use a single A100 GPU for the 1.5B, 7B and 8B models, and 4 A100 GPUs for the 70B models to perform the search. We make use of device_map in huggingface for multi-GPU execution, which is sequential execution of layers on different devices. The 7B/8B models require about 7 minutes and the 70B models require less than 60 minutes to complete the search. Considering that the quantization only needs to be performed once, the searching algorithm is practical for the real workloads.
5 Summary
We have presented MixLLM, achieving high accuracy with low memory consumption and high system efficiency with the rarely explored optimization space of mixed-precision quantization between output features. MixLLM identifies the salience of each output feature according to the loss distance estimation w.r.t. the global model loss rather than local layer loss. By assigning larger bit-width to the features need it most, MixLLM achieves the superior accuracy to SOTA with low memory consumption. The sub-problems of different bit-widths are disjoint and can run in parallel efficiently on the GPU. We have identified the sweet spot of the quantizaiton configuration that is friendly to both accuracy and system efficiency. To address the challenge of system efficiency, we design the two-step dequantization to enable using int8 Tensor Core computation and the fast integer-float conversion to reduce the dequantization overhead. We have designed the end-to-end software pipeline to overlap the memory access, the dequantization computation with SIMT Core and the MatMul with Tensor Core. Experiment results show that MixLLM achieves superior accuracy to existing works and state-of-the-art system efficiency with low memory cost.

### v2

@@
-MixLLM: LLM Quantization with Global Mixed-precision between Output-features and Highly-efficient System Design
+
 Abstract
 Quantization has become one of the most effective methodologies to compress LLMs into smaller size.
+However, the existing quantization solutions still show limitations of either non-negligible accuracy drop or low system efficiency.
+In this paper, we propose MixLLM that explores the optimization space of mixed-precision quantization between output features, based on the insight that different features matter differently in the model.
+MixLLM identifies the important output features in the global view rather than within each single layer, effectively assigning larger bit-width to output features that need it the most to achieve high accuracy and low memory usage.
+We present the sweet spot of quantization configuration of algorithm-system co-design with high accuracy and system efficiency.
+To address the system challenge, we design the two-step dequantization to make use of the Tensor Core easily and fast data type conversion to reduce dequantization overhead, and present the software pipeline to overlap the memory access, dequantization and the MatMul to the best.
+Extensive experiments show that with only 10% more bits, the perplexity increase can be reduced from about 0.5 in SOTA to within 0.2 for Llama 3.1 70B, while MMLU-Pro loss can be reduced from 1.92 to 0.99 over the SOTA of three popular models.
+Besides its superior accuracy, MixLLM also achieves state-of-the-art system efficiency.
+Code is released at https://github.com/microsoft/MixLLM.
+footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
+Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
+1 Introduction
+Large language models (LLMs) Bubeck et al.
+(2023); Meta (Cited 2024) have shown remarkable performance on various tasks.
+But their large memory consumption and massive computation cost have become an obstacle for the efficient deployment Xia et al.
+(2023; 2024).
+Quantization has become one of the most effective solution to compress LLMs into smaller size Frantar et al.
+(2022); Lin et al.
+(2024); Xiao et al.
+(2023); Yao et al.
+(2022), by representing the weight or activation with smaller bit-width.
 However, the existing quantization solutions still show limitations of either non-negligible accuracy drop or system inefficiency.
-In this paper, we make a comprehensive analysis of the general quantization principles on their effect to the triangle of accuracy, memory consumption and system efficiency.
-We propose MixLLM that explores the new optimization space of mixed-precision quantization between output feature based on the insight that different output feature matter differently in the model.
-MixLLM identifies the output feature with high salience in the global view rather than within each single layer, effectively assigning the larger bit-width to output features that need it most to achieve good accuracy with low memory consumption.
-We present the sweet spot of quantization configuration of algorithm-system co-design that leads to high accuracy and system efficiency.
-To address the system challenge, we design the two-step dequantization to make use of the int8 Tensor Core easily and fast data type conversion to reduce dequantization overhead significantly, and present the software pipeline to overlap the memory access, dequantization and the MatMul to the best.
-Extensive experiments show that with only 10% more bits, the PPL increasement can be reduced from about 0.5 in SOTA to within 0.2 for Llama 3.1 70B, while on average MMLU-Pro improves by 0.93 over the SOTA of three popular models.
-In addition to its superior accuracy, MixLLM also achieves state-of-the-art system efficiency.
-1 Introduction
-Large language models (LLMs) [7, 32] have shown remarkable performance on various tasks.
-But their large memory consumption and massive computation cost have become an obstacle for the efficient deployment [43, 44].
-Quantization has become one of the most sufficient solution to compress LLMs into smaller size [14, 26, 45, 46], by representing the weight or activation with smaller bit-width.
-However, the existing quantization solutions still show limitations of either non-negligible accuracy drop or system inefficiency.
-There is a triangle of characteristics for efficient LLM quantization: accuracy, memory consumption of parameters, and system efficiency of execution, which we call effectiveness triangle of quantization.
-The existing quantization solutions have different focus and trade-off in the triangle:
-  • The weight-only methodologies target to solve the memory consumption problem, and can speedup the small-batched decoding execution that faces the memory-wall problem [43, 21].
-But their accuracy drop of 4-bit quantization can be a challenge for the production workloads sensitive to accuracy, which becomes more serious in the new models with higher information density like Llama 3 [32], as illustrated in recent studies [42, 22].
-Besides, the weight-only method can lead to system performance drop for large-batched workloads (e.g., the SOTA W4A16 kernel only achieves 83% performance of its float16 counterpart at batch size 512 with hidden size 4096, shown in Fig.5).
-  • The weight-activation quantization represents the activation with low-bit values along with the weights, potentially lead to higher system efficiency.
-But it can lead to larger accuracy drop than the weight-only method as the activation is usually harder to quantize [50, 3, 27].
+There is a triangle of characteristics for LLM quantization: accuracy, memory consumption of parameters, and system efficiency of execution, The existing quantization solutions have different focus and trade-off in the triangle: 1) The weight-only method targets to solve the memory consumption problem, and can speedup the small-batched decoding execution that faces the memory-wall problem Xia et al.
+(2023); Kim et al.
+(2024).
+But their accuracy drop of 4-bit quantization can be a challenge for the production workloads sensitive to accuracy, as illustrated in recent studies Wu et al.
+(2023); Kumar et al.
+(2024).
+Besides, the weight-only method can lead to system performance drop for large-batched workloads due to the dequantization overhead.
+2) The weight-activation quantization represents the activation with low-bit values along with the weights, potentially lead to higher system efficiency.
+But it can lead to larger accuracy drop than the weight-only method as the activation is usually harder to quantize Zhao et al.
+(2024); Ashkboos et al.
+(2024); Lin et al.
+(2025).
 Besides, it introduces more dequantization overhead for the activation that can hurt the system efficiency.
-The transformation optimizations in some works can make the system efficiency even worse.
-  • Outlier separation and mixed-precision technologies emerge to improve the accuracy of low-bit quantization by either excluding the unstructured high-salience weights from quantization [11, 21] or assigning larger bit-width for the quantization of structured high-salience weights [50].
-The former shows system efficiency problem due to the low efficiency of half precision (i.e., float16/bfloat16) sparse tensor processing.
-The state-of-the-art mixed-precision solution [50] aims for low-bit quantization but shows non-negligible accuracy drop, even inferior to the 4-bit weight-only quantization.
 Figure 1: Illustration of the quantization with mixed-precision between output features and kernel execution.
-Contributions.
-In this paper, we provide an extensive analysis of the general quantization principles.
-To address the limitations of the previous works and cover the three characteristics in the effectiveness triangle, we propose MixLLM, which makes the following contributions:
+Contributions. In this paper, we provide an extensive analysis of the general quantization principles, and propose MixLLM with the following contributions:
 […]
 […]
-Extensive evaluation shows that, with only 10% of 8-bits (i.e., W4.4A8), MixLLM outperforms all the existing 4-bit quantization algorithms while achieving the state-of-the-art system efficiency.
+Extensive evaluation shows that, with 10% 8-bit and 90% 4-bit (i.e., W4.4A8), MixLLM significantly outperforms SOTA quantization algorithms while achieving the state-of-the-art system efficiency on both A100 and H100 GPUs.
 2 Background, Related Work, and Discussion
-2.1 Background of Quantization
-The quantizaiton maps the tensor $X$ into the target range with smaller bit-width representation through affine transformation: $X_{q}=clamp(\lfloor\frac{X}{s}\rceil+z,range)$, where $s$ is the scale and $z$ is the zero point.
-The value can be recovered (i.e., dequantization) through: $X^{\prime}=(X_{q}-z)\times s$.
-$X^{\prime}$ is pushed to the discrete chunks rather than recovered to the original value, thus has accuracy loss.
-The bit-width is essential for the accuracy of quantization as it determines the number of chunks for the quantized values ($2^{bit\_width}$).
-Take an example, enlarging the bit-width from 4 to 5 can double the number of chunks, so that the 5-bit RTN quantization can easily beat the 4-bit quantizations of advanced techniques (Tab.1).
-The scale and zero point can be calculated from the whole channel/token vector or a small group within the channel/token, the former is called per-channel/token quantization and the latter is group-wise quantization.
-The group-wise scheme results in smaller accuracy loss due to the smaller chunk scale, but requires more complex GPU kernel design.
-The symmetric quantization uses 0 as the zero point value, which simplifies the computations ($X_{q}=clamp(\lfloor\frac{W}{s}\rceil,range)$, $X^{\prime}=X_{q}\times s$) and enables many works to design the per-channel/per-token quantized kernels by multiplying the scales at the epilogue of the whole MatMul (matrix multiplication) for dequantization [45, 40].
-However, it leads to larger loss than the asymmetric one as the data distribution can be usually asymmetric, especially for smaller bit-widths like 4-bit.
-2.2 Related Works and Discussion of General Quantization Principles
-This paper mainly focuses on post-training quantization (PTQ).
+Quantization maps the tensor $X$ into the target range with smaller bit-width through $X_{q}=clamp(\lfloor\frac{X}{s}\rceil+z,range)$, where $s$ is the scale and $z$ is the zero point.
+The scale and zero point can be calculated from the whole channel/token vector or a smaller group.
+The group-wise scheme results in higher accuracy but requires more complex GPU kernel design.
+The symmetric quantization uses 0 as the zero point, which simplifies the computations and enables many works to design the per-channel/per-token quantized kernels by multiplying the scales at the epilogue of the whole MatMul (matrix multiplication) for dequantization Xiao et al.
+(2023); TensorRT-LLM (Cited 2024).
+However, it usually leads to larger loss than the asymmetric one, especially for smaller bit-widths like 4-bit.
+2.1 Related Work and Discussion
+This paper focuses on pure post-training quantization (PTQ), without the training on the weight (i.e., QAT) or the quantization parameters (e.g., training the rotation matrix like in SpinQuant Liu et al.
+(2024)).
 Systems that affect the quantization requirement.
-The continuous batching technology [47] enables to batch the decoding tasks from different requests together to enlarge the batch dimension of MatMul during LLM inference.
-The chunked-prefill method [19, 2, 51] advances the continuous batching by merging the prefill and decoding tasks into the same batch, further enlarging the MatMul shapes.
+The continuous batching technology Yu et al.
+(2022) enables to batch the decoding tasks from different requests together to enlarge the batch dimension of MatMul during LLM inference.
+The chunked-prefill method Holmes et al.
+(2024); Agrawal et al.
+(2024); Zheng et al.
+(2024) advances the continuous batching by merging the prefill and decoding tasks into the same batch, further enlarging the MatMul shapes.
 These technologies pushes many LLM jobs to become compute-bound and motivate the demand to reduce computation.
@@
 There emerges a wide range of technologies to improve the accuracy of weight-only quantization.
-GPTQ [14] advances OBC [13] on OBS-based [18] weight compensation with blocked updating and reordering.
-AWQ [26] proposes to scale the weight according to the characteristic of activation.
-OminiQuant [37]) proposes the learnable scaling and weight clipping factors.
-SpQR [11], SqueezeLLM [21] and OWQ [25] separate the outliers from the quantiation and with half precision.
-QuiP [8] aims to achieve extreme low-bit quantization with incoherence processing.
-ZeroQuant(4+2) [42] aims to improve accuracy with medium-sized FP6 quantization.
-The weight-only quantization does not reduce the computation but introduces the extra dequantization operations.
-The low-bit weight will be dequantized to float16 to execute the MatMul in float16 datatype.
-The current weight-only quantization faces two challenges: 1) From the accuracy aspect, there is still an accuracy gap between the 4-bit quantization and the float16 model, especially for many real business scenarios sensitive to the small accuracy drop, as discussed in the recent works [42, 44].
+GPTQ Frantar et al.
+(2022) advances OBC Frantar and Alistarh (2022) on OBS-based Hassibi et al.
+(1993) weight compensation with blocked updating and reordering.
+AWQ Lin et al.
+(2024) proposes to scale the weight according to the characteristic of activation.
+OmniQuant Shao et al.
+(2024) proposes the learnable scaling and weight clipping factors.
+SpQR Dettmers et al.
+(2024), SqueezeLLM Kim et al.
+(2024) and OWQ Lee et al.
+(2024) separate the outliers from the quantiation and with half precision.
+QuIP Chee et al.
+(2023) and QuIP# Tseng et al.
+(2024a) aim to achieve extremely low-bit quantization with incoherence processing, QTIP Tseng et al.
+(2024b) also leverages incoherence processing.
+AQLM Egiazarian et al.
+(2024) and PV-Tuning Malinovskii et al.
+(2024) target extremely low-bit quantization.
+SliM-LLM Huang et al.
+(2024) focuses on extremely low-bit quantization using intra-layer mixed-precision.
+ZeroQuant(4+2) Wu et al.
+(2023) and Quant-LLM Xia et al.
+(2024) aim to improve accuracy with FP6 quantization.
+The weight-only quantization does not reduce the computation but introduces the extra dequantization operations, as it requires to dequantize the weight into float16 for execution.
+The current weight-only quantization faces two challenges: 1) From the accuracy aspect, there is still an accuracy gap between the 4-bit quantization and the float16 model, especially for many real business scenarios sensitive to the small accuracy drop, as discussed in the recent works Wu et al.
+(2023); Xia et al.
+(2024).
 2) It can lead to system efficiency drop on busy servers as the recent LLM inference serving systems will usually batch the processing of different requests together on the server and form large MatMuls.
-The large MatMuls are compute-bound and will suffer from the dequantization overhead [27].
+The large MatMuls are compute-bound and will suffer from the dequantization overhead Lin et al.
+(2025).
 Weight-activation quantization and the challenges.
 The weight-activation quantization helps to make use of the low-bit computing unit.
-LLM.int8() [10] observes the activation outlier problem and separates outliers from quantization with half precision.
-ZeroQuant [46] proposes the per-token activation quantization and group-wise weight quantization.
-SmoothQuant [45] addresses the activation outlier problem through smoothing, and AffineQuant [29] proposes the general affine transformation for quantization.
-RPTQ [48] reorders the channels to cluster similar scaled values together.
-SpinQuant [28] and QuaRot [3] leverages matrix rotation properties to alleviate the outlier phenomenon.
-Atom [50] uses the mixed-precision between input features to improve accuracy of 4-bit activation quantization.
-QoQ [27] is a holistic quantization solution with progressive group quantization, attention smoothing, rotation, and channel reordering.
-Even though the weight-activation quantization has the advantage of reduced MatMul computation (i.e., MatMul in smaller bit-width to make use of the smaller bit-width computing unit with higher computation throughput [^2] The extra dynamic activation quantization kernel can be fused into other operators with very little system cost [50], thus we only discuss the MatMul itself.), it faces the challenge of accuracy drop caused by the activation quantization, especially that the activation is usually harder to quantize than the weight.
-The SOTA low-bit weight-activation solutions [3, 28, 27] still have a gap to the 4-bit weight only quantization.
-Beside the accuracy drop, the activation quantization will introduce more dequantization overhead than the weight-only one, which makes it challenging to design efficient GPU kernels.
+LLM.int8() Dettmers et al.
+(2022) observes the activation outlier problem and separates outliers from quantization with half precision.
+ZeroQuant Yao et al.
+(2022) proposes the per-token activation quantization and group-wise weight quantization.
+SmoothQuant Xiao et al.
+(2023) addresses the activation outlier problem through smoothing, and AffineQuant Ma et al.
+(2024) proposes the general affine transformation for quantization.
+RPTQ Yuan et al.
+(2023) reorders the channels to cluster similar scaled values together.
+SpinQuant Liu et al.
+(2024) and QuaRot Ashkboos et al.
+(2024) leverages matrix rotation properties to alleviate the outlier phenomenon.
+Atom Zhao et al.
+(2024) uses the mixed-precision between input features to improve accuracy of 4-bit activation quantization.
+QServe Lin et al.
+(2025) is a holistic solution of W4A8 quantization.
+Even though the weight-activation quantization has the advantage of reduced MatMul computation (i.e., smaller bit-width computation with higher throughput), it faces the challenge of accuracy drop caused by the activation quantization.
+The SOTA low-bit weight-activation solutions Ashkboos et al.
+(2024); Liu et al.
+(2024); Lin et al.
+(2025) still have a gap to the 4-bit weight only quantization.
+Besides the accuracy drop, the activation quantization will introduce more dequantization overhead than the weight-only one, which makes it challenging to design efficient GPU kernels.
 When enabling the asymmetric quantization, the result of $(X_{q}-z)$ may exceed the range of the bit-width of $X_{q}$, making it hard to use the corresponding Tensor Core computation.
-Systems like Atom [50] thus avoid using the asymmetric quantization, with the cost of larger accuracy drop.
-The group-wise quantization requires fine-grained integer-to-float (I2F) conversion to apply per-group scales.
-However, the I2F instruction is more expensive than the common computation instructions on the GPU [1] and can lead to large system performance drop ($>10\%$ drop in our practice).
-Besides, the throughput of Tensor Core is much higher than that of SIMT Cores, 624 TOPS of int8 Tensor Core vs.
-19.5 TFLOPS/TOPS of FP32/INT32 SIMT Cores.
-There still lacks a well designed software pipeline to overlap the Tensor Core computation and SIMT Core based dequantization in the existing works while achieving a high accuracy.
-In general, the existing solutions focus on partial of the effectiveness triangle, but cannot cover all of them well.
-MixLLM is orthogonal to the above works by exploring the mixed-precision between output features with global salience identification, and the co-designed quantization decision and GPU kernels.
+Systems like Atom Zhao et al.
+(2024) thus avoid using the asymmetric quantization, with the cost of larger accuracy drop.
+The integer quantization requires integer-to-float (I2F) conversion to apply scales.
+However, the I2F instruction is more expensive than the common computations on modern GPUs Abdelkhalik et al.
+(2022) and can lead to large system performance drop for group-wise quantization ($10\%$ drop in our practice).
+Besides, the throughput of Tensor Core is much higher than SIMT Cores, 624 TOPS of int8 Tensor Core vs.
+19.5 TFLOPS/TOPS of FP32/INT32 SIMT Cores on A100 GPU.
+There lacks a well designed software pipeline to overlap the Tensor Core computation and SIMT Core based dequantization for the group-wise and asymmetric weight-activation quantization.
+Performance challenge of the float16 outlier separation Outlier separation with half precision works to improve the accuracy while using small bit-width for the non-sensitive weights Kim et al.
+(2024); Dettmers et al.
+(2024), by separating the outliers into an extra sparse tensor in float16.
+However, it is hard to achieve the peak performance due to the inefficiency of the sparse computation on the GPU, especially when the batch size is large and the linear layer becomes compute-bounded.
+(As discussed in Flash-LLM Xia et al.
+(2023), the hardware utilization can be lower than 10% for the sparse MatMul, while its dense counterpart can usually achieve more than 60%.) This is because the unstructured tensor computation cannot make use the fast Tensor Core easily, but has to use the SIMT Core in float16 for computation and float32 for accumulation [^1] Flash-LLM Xia et al.
+(2023) optimizes the unstructured sparse MatMul, but can only speedup the small-batched scenarios..
+Moreover, sparse computing makes it more difficult to fully utilize the hardware due to the non-continuous memory pattern and the extra index computation.
 3 Methodology
 3.1 Quantization Design and Decision in MixLLM
-To cover the three aspects of the effectiveness triangle simultaneously, we make the following design and decision of weight and activation quantization according to the analysis in Sec.2.2.
-3.1.1 Mixed-precision between output features of weight, with global salience identification.
-It is known that different elements of the weight show different salience to the network’s loss when being quantized [21, 11].
-The outlier separation method can improve the accuracy by using float16 to store the high-salience elements, but can suffer from the inefficient sparse MatMul.
+We make the following design and decision to optimize the quantization algorithm and system.
+3.1.1 Global Mixed-precision
+Different elements of the weight show different salience to the network’s loss when being quantized Kim et al.
+(2024); Dettmers et al.
+(2024).
+The outlier separation method can improve the accuracy by using float16 for high-salience elements, but can suffer from the inefficient sparse MatMul.
 We observe that the elements with high salience tend to show distribution along the output channels for most of the linear layers in many LLMs.
-Based on this observation, we can assign larger bit-width to the output channels of high salience, and smaller bit-with to the others, forming structured mixed-precision quantization.
-Through the experiments, we get the same conclusion with the existing works [21, 11] that there is only a small set of elements with high salience contributing significantly to the model’s accuracy drop.
-Thus we only need to assign the large bit-width to a small portion of the output channels to achieve good accuracy and retain a small memory consumption at the same time.
-The structured mixed-precision between different output channels can be friendly to the system efficiency and kernel development, due to the nature that different output features are disjoint in the MatMul and the computation of them are different sub-problems.
-Fig.1 shows how the linear layer computes with the mixed-precision between output features.
-It divides the linear into independent sub-problems, and finally gathers the output of the sub-problems together to form the final result.
-This optimization space is orthogonal to the existing quantization optimizations, e.g., GPTQ [14], and can be applied together with them.
+Based on this observation, we can assign larger bit-width to the output channels of high salience, and smaller bit-width to the others.
+Through the experiments, we get the same conclusion with the existing works Kim et al.
+(2024); Dettmers et al.
+(2024) that there is only a small set of elements with high salience to quantization.
+Thus we only need to assign the large bit-width to a small portion of the output channels to achieve good accuracy while retaining a small memory consumption.
+The structured mixed-precision between output channels can be friendly to the system efficiency and kernel development, due to the nature that different output features are disjoint in the MatMul and the computation of them are different sub-problems.
+Fig.1 illustrates the mixed-precision quantization and execution in MixLLM.
+It divides the linear into independent sub-problems, and finally gathers the output of the sub-problems together to form the result.
+This optimization space is orthogonal to the existing quantization methodologies and can be applied together with them.
+Figure 2: The percentage of high-salient out features within each linear layer of Llama 3.1 8B model according to each feature’s contribution to the final loss after quantizing to 4-bit, with 10% high-salient features globally.
+Each decoder layer contains q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj in order.
 One critical problem is how to identify the high-salience output channels in the model.
-The fixed threshold [11] or the fixed number/ratio [50, 25] of high salience elements computed by the local loss of layers can be sub-optimal to the end-to-end model, as different layers can show different importance to the model’s final output [16, 30, 12].
+The fixed threshold Dettmers et al.
+(2024) or the fixed number/ratio Zhao et al.
+(2024); Lee et al.
+(2024) of high salience elements computed by the local loss of layers can be sub-optimal to the end-to-end model, as different layers show different importance to the model’s final output Gromov et al.
+(2024); Men et al.
+(2024); Dong et al.
+(2019).
 A high salience channel w.r.t.
@@
 As a result, different layers will have different number of high salience channels.
-Fig.2 shows the distribution of the top 10% high-salient out features in Llama 3.1 8B, showing huge difference in different linear layers.
-Figure 2: The percentage of high-salient out features within each linear layer of Llama 3.1 8B model according to each feature’s contribution to the final loss after quantizing to 4-bit, with 10% high-salient features globally.
-Each decoder layer contains q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj in order.
-Note that this design is different from the mixed-precision in Atom [50] from two aspects.
+Fig.2 shows the distribution of the top 10% high-salient out features in Llama 3.1 8B, showing huge difference between different layers.
+Note that this design is different from the mixed-precision in Atom Zhao et al.
+(2024) from two aspects.
 1) MixLLM first addresses the problem of identifying the high-salience channels globally rather than locally.
-2) MixLLM applies the mixed-precision between output features rather than input features, which is more system performant and algorithm flexible [^3] One example is that, the outlier number in each MatMul should be a multiplier of the corresponding tiling size of kernel design to achieve a good system efficiency in Atom, which limits the flexibility of algorithm.
-as the output features are disjoint naturally.
-3.1.2 Quantization decision with algorithm-system consideration: 8-bit symmetric activation and 4-bit asymmetric weight quantization in group-wise manner.
-MixLLM makes the same decision with QoQ [27] on activation quantization to use 8-bit, as the 4-bit activation can lead to a large accuracy drop but does not lead to significant system efficiency improvement as MatMul execution tends to be bound more on the larger weight tensor rather than the smaller activation tensor.
-It can be partially indicated from the compute intensity of the linear layer.
-Given token number $M$ and in/out features $K$/$N$, the compute intensity $I=\frac{2MNK}{MKB_{act}+KNB{weight}}$.
-$B_{act}$ and $B_{weight}$ are the bytes per element of activation and weight.
-Given $M=512$ and $N=K=4096$, reducing $B_{weight}$ from 8 to 4 will results in an 80% increasement of $I$, while reducing $B_{act}$ from 8 to 4 will only achieve 5.88% increasement.
+2) MixLLM applies the mixed-precision between output features rather than input features, which is more system/algorithm flexible as the output features are disjoint naturally.
+It also differs from SliM-LLM Huang et al.
+(2024) which also only considers the local loss to determine the mixed precision and does not focus on the system performance problem.
+3.1.2 Quantization Decision
+MixLLM makes the same decision with QServe Lin et al.
+(2025) on activation quantization to use 8-bit, as the 4-bit activation can lead to a large accuracy drop but does not lead to significant system efficiency improvement as MatMul execution tends to be bound more on the larger weight tensor rather than the smaller activation tensor according to the compute intensity $I=\frac{2MNK}{MKB_{act}+KNB{weight}}$ ($M$ is the token number, KK/NN are the in/out features, and $B_{act}$ and $B_{weight}$ are the bytes per element of activation and weight).
 Instead of using per-token activation quantization, MixLLM uses group-wise RTN method.
-On the one hand, Tab.1 shows that simple group-wise RTN quantization performs better than token-wise smoothing method.
-On the other hand, the weight is already group-wise in MixLLM, and the group-wise activation does not lead to significant more dequantization overhead in the system.
-We observe symmetric quantization is enough for the 8-bit activation (refer to MixLLM W8A8 in Tab.1), while asymmetric can be essential for the 4-bit weight.
-The group-wise method with asymmetric can lead to a difficulty for the kernel to make use int8 Tensor Core, for which QoQ [27] introduces the two-step quantization method.
-Instead, we design a two-step dequantization method with the property of the mix of symmetric and asymmetric (Sec.3.3).
+On the one hand, Tab.1 shows that simple group-wise RTN quantization outperforms token-wise smoothing method.
+On the other hand, the weight is already group-wise in MixLLM, and the group-wise activation does not lead to significantly more dequantization overhead in the system.
+We observe symmetric quantization is enough for the 8-bit activation (refer to MixLLM W8A8 in Tab.1), while asymmetric is essential for the 4-bit weight.
+The group-wise method with asymmetric can lead to difficulty for the kernel to make use int8 Tensor Core, for which we design a two-step dequantization method with the property of the mix of symmetric and asymmetric (Sec.3.3).
 3.2 Global Precision Search Algorithm
-As discussed in Sec.3.1, MixLLM determines the precision of all output features in all layers globally rather than locally.
+MixLLM determines the precision of all output features in all layers globally.
 It identifies the salience of these features with respect to the final loss of the model, and assigns larger bit-width to the features leading to larger loss.
@@
 $S_{c}=\lvert l(c_{q})-l(c_{0})\rvert$ (1)
-which is the distance of the model’s loss between quantizing and not quantizing this single channel.
+which is the distance of the model’s loss between quantizing and not quantizing channel $c$.
 In Eq.1, $l()$ is the loss function of the model w.r.t.
-a single channel, $c_{q}$ is the quantized weight of the channel and $c_{0}$ is the original weight.
-Note that it regards other neurons except $c$ as constant in $l()$.
-We use the Taylor Expansion method to estimate the loss function $l(c)$ (similar with the existing quantization works, ignoring the high-order items):
+a single channel, $c_{q}$ is the quantized weight of the channel and $c_{0}$ is the original one.
+We use the Taylor Expansion method to estimate the loss function $l(c)$, ignoring the high-order items:
 $l(c)\approx l(c_{0})+g^{T}(c-c_{0})+\frac{1}{2}(c-c_{0})^{T}H(c-c_{0})$ (2)
 where $g=\mathbb{E}[\frac{\partial}{\partial c}l(c)]$ is the loss’s gradient w.r.t.
-the channel, and $H=\mathbb{E}[\frac{\partial^{2}}{\partial c^{2}}l(c)]$ is the second-order gradient (i.e., Hessian matrix) w.r.t.
-the channel.
-It is infeasible to calculate the Hessian matrix as it is too costly.
-We approximate the Hessian $H$ with the (empirical) Fisher information matrix $\mathit{F}$ on the calibration dataset $\mathit{D}$:
-$H\approx\mathit{F}=\frac{1}{\lvert\mathit{D}\rvert}\sum_{d\in\mathit{D}}{g_{d}g_{d}^{T}}$ (3)
-Note that $\mathit{F}$ is w.r.t. a channel, differing from the diagonal Fisher information matrix in the recent works that ignores any cross-neuron interactions [23, 21].
+the channel, and $H=\mathbb{E}[\frac{\partial^{2}}{\partial c^{2}}l(c)]$ is the second-order gradient (i.e., Hessian matrix).
+It is infeasible to calculate the Hessian matrix.
+We approximate the Hessian with the (empirical) Fisher information matrix (FIM) $F$ on the calibration dataset $\mathcal{D}$:
+$H\approx F=\frac{1}{\lvert\mathcal{D}\rvert}\sum_{d\in\mathcal{D}}{g_{d}g_{d}^{T}}$ (3)
+Note that $F$ is w.r.t. a channel, differing from the diagonal FIM in the recent works that ignores any cross-neuron interactions Kwon et al. (2022); Kim et al. (2024).
 Based on this approximation, the second order loss factor $\frac{1}{2}(c-c_{0})^{T}(g_{d}g_{d}^{T})(c-c_{0})$ can be further simplified to $\frac{1}{2}(g_{d}^{T}(c-c_{0}))^{2}$, simplifying the expensive chained matrix multiplication into a single vector product.
@@
 $S_{c}=\frac{1}{\lvert\mathit{D}\rvert}\sum_{d\in\mathit{D}}\lvert g_{d}^{T}(c_{q}-c_{0})+\frac{1}{2}(g_{d}^{T}(c_{q}-c_{0}))^{2}\rvert$ (4)
-We do not ignore the first order information during the calculation, differing from OBD [24] and many recent quantization works [14, 11, 21].
-This is because the first order factor can be more significant in the estimation in Eq.4, as the estimated second order factor is the square of the first order factor divided by two for each sample.
-Considering that $g$ can be very small for the well pretrained models and the delta of the quantized weight is usually not large, the first order factor can be larger than the second order one.
-Besides, what we require is the loss itself rather than the arguments of the loss function, and thus we do not need to ignore the first order factor to simplify the arguments calculation.
+We do not ignore the first order information during the calculation, differing from the recent quantization works Frantar et al.
+(2022); Dettmers et al.
+(2024); Kim et al.
+(2024).
+Note that what we require is the loss itself rather than the arguments of the loss function, thus we do not need to ignore the first order factor to simplify the arguments calculation.
 Algorithm 1 Global precision search procedure.
-1:Linear layer number $L$, weight and gradient of all linear layers ($W_{i}\in\mathbb{R}^{O,I}$, $G_{i}\in\mathbb{R}^{O,I}$ for layer $i$).
-2:Global channel index with large and small bit width (largebit_channels, smallbit_channels).
-3:$S_{global}$leftarrow []
-4:for $i=1,2,...,L$ do
-5:     $W_{i}^{delta}$leftarrow quantize($W_{i}$) - $W_{i}$
-6:     $S_{1st}$leftarrow sum($G_{i}\odot W_{i}^{delta}$, dim=1) $\triangleright$ Per-channel dot product between $G_{i}$ and $W_{i}^{delta}$
-7:     $S_{2nd}\leftarrow 0.5*(S_{1st})^{2}$
-8:     $S\leftarrow\lvert S_{1st}+S_{2nd}\rvert$triangleright $S\in\mathbb{R}^{O}$, the salience of the $O$ channels
-9:     for $channel\_id=1,2,..,O$ do $\triangleright$ Log the salience of each output channel of this layer
-10:         $S_{global}$.append(tuple($i$, $channel\_id$, $S[channel\_id]$))
-11:sort($S_{global}$) $\triangleright$ Sort according to the salience, descending
-12:largebit_channels, smallbit_channels $\leftarrow$ $S_{global}[:N_{largebit}]$, $S_{global}[N_{largebit}:]$
+0: Weight and gradient of all linear layers ($W_{i}\in\mathbb{R}^{O\times I}$, $G_{i}\in\mathbb{R}^{O\times I}$ for layer $i\in[1..L]$). Total number of output channels with large bit-width ($T$).
+0: Channel index for large and small bit-width ($\mathcal{C}^{lg}$ and $\mathcal{C}^{sm}$).
+[…]
+2:for $i=1$ to $L$ do
+3:  $W_{i}^{delta}$ $\leftarrow$ quantize($W_{i}$) - $W_{i}$
+4:  $S^{1st}$ $\leftarrow$ sum($G_{i}\odot W_{i}^{delta}$, dim=1)
+5:  $S^{2nd}\leftarrow 0.5*(S^{1st})^{2}$
+6:  $S\leftarrow\lvert S^{1st}+S^{2nd}\rvert \triangleright$ $O$ channels’ salience $\in\mathbb{R}^{O}$
+7:  for $cid=1$ to $O$ do
+8:   $\mathcal{C}^{global}$ $\leftarrow$ $\mathcal{C}^{global}$ $\cup$ ( tuple(ii, cidcid, $S_{cid}$) )
+9:  end for
+10:end for
+11: Sort $\mathcal{C}^{global}$ by salience in descending order.
+[…]
 Algo.1 illustrates the procedure of the global precision search.
-It calculates the salience of all the output channels of all linear layers and sort them in descending order globally.
-Given the global threshold $N_{largebit}$ as the number of large-bit precision channels, the first $N_{largebit}$ channels are intended to be quantized with 8-bit, and the other channels will be quantized with smaller bit-width (i.e., 4-bit in this paper).
-Any quantization methodologies (e.g., GPTQ, clip search) can be applied independently to these two disjoint parts of channels.
-Note that we calculate the salience of the channels in one pass rather than iterative identifying the high-salience parts in a smaller step, as we observe the single-pass method show similar results with the iterative method and saves a lot of computation overhead than the latter.
+It calculates the salience of all the output channels of all linear layers and sort them in descending order.
+Given the global threshold $T$ as the number of large-bit precision channels, the first $T$ channels are intended to be quantized with 8-bit, and the other channels will be quantized with smaller bit-width (i.e., 4-bit in this paper).
 3.3 Efficient Quantization Computation System
+Figure 3: The GPU kernel software pipeline of group-wise W4A8/W8A8 quantized MatMul.
+It assumes perfect overlapping.
+G2S: load global to shared memory; S2R: load shared memory to register; MMA: matrix multiply-accumulation; I2F: integer to float conversion; deq: dequantize; acc: accumulate.
+While this pipeline is modeled on the NVIDIA A100 architecture, its fundamental principles remain applicable to subsequent generations, such as Hopper and Blackwell, subject to minor architectural refinements.
+For instance, newer architectures can utilize the Tensor Memory Accelerator (TMA) to load activation tensors directly, bypassing registers before they reach the Tensor Cores.
+Furthermore, these architectures support warp specialization for memory loading as an alternative to a uniform execution scheme.
 Two-step dequantization to make use of int8 Tensor Core.
-As for the W4A8 computation, the dequantized weight and activations are $(W_{q}-z)s_{w}$ and $A_{q}s_{a}$, where $W_{q}$ and $z$ are uint4 datatype, $A_{q}$ is int8 datatype, and $s_{w}$ and $s_{a}$ are float16 datatype.
+As for the W4A8 computation, the dequantized weight and activations are $(W_{q}-z)s_{w}$ and $A_{q}s_{a}$, where $W_{q}$ and $z$ are uint4 datatype (4-bit unsigned integer), $A_{q}$ is int8 datatype, and $s_{w}$ and $s_{a}$ are float16 datatype.
 Directly dequantizing the tensors into float16 datatype before the MatMul computation will prevent us using the fast 8-bit Tensor Core on the GPU.
@@
 Figure 4: The float and integer value of binary (010010110xx...x), each within a consecutive range.
-Fast Int to Float conversion with partially fusing into Tensor Core instruction.
-In the above two-step dequantization computation, the step 2 is the MatMul between the integer tensor $A_{q}(W_{q}-z)$ and the float tensor $s_{a}s_{w}$.
-It requires the integer to float conversion (I2F) before the multiply operation.
-The I2F instruction is expensive on the modern GPUs.
-Instead, we make use of the range-dependent fast I2F transformation to convert the I2F instruction into two add/sub instructions [^4] CUTLASS [9] also has an implementation of fast I2F for general purpose..
+Fast I2F with partially fusing into Tensor Core instruction.
+In the above two-step dequantization computation, the step 2 is the multiplication between integer and float tensor, requiring integer to float conversion (I2F).
+As I2F instruction is expensive on the modern GPUs, we make use of the range-dependent fast I2F transformation to convert the I2F instruction into two add/sub instructions.
 Specifically, it is based on the fact that there exists a certain range where an integer value’s binary is the same as a corresponding float binary.
 As shown in Fig.4, the binary with the first 9 bits as 010010110 represents a series of consecutive int32 and float32 values, respectively.
-We can add a bias to an integer value to make it within this consecutive range, and then subtract a corresponding bias in float (same underling binary) to restore its value in float type.
-We take the middle value in this range as the bias to maximize the data range that can be safely converted, whose hexadecimal number is 0x4b400000 (i.e., in the remaining 23 bits, the first bit is 1 and the other bits are 0).
+We can add a bias to an integer value to make it within this consecutive range, and then subtract a corresponding bias in float to restore its value in float.
+We take the middle value in this range as the bias to maximize the data range that can be safely converted, whose hexadecimal number is $mid=\texttt{0x4b400000}$ (i.e., in the remaining 23 bits, the first bit is 1 and the other bits are 0).
 This allows to convert a consecutive range of $2^{23}$ int32 numbers to float32.
-The range of dot product of $k$ int8 elements is $2^{16}k$, thus the above fast I2F conversion allows the $k$ value up to 128.
+The range of dot product of $k$ int8 elements is in $2^{16}k$, thus the above fast I2F conversion allows the $k$ value to be 128.
 We use quantization group size as 128 and can use the fast I2F safely:
-1// bias_int = as_int(0x4b400000), bias_fp = as_float(0x4b400000);
+1// bias_int = as_int(mid), bias_fp = as_float(mid);
 2int tmp = src_int + bias_int;
 3int dst_float = *((float*)&tmp) - bias_fp;
-We further fuse the integer subtraction into the Tensor Core mma (Matrix Multiply-Accumulate) instruction.
-The mma instruction computes $D=AB+D$ during the MatMul computation.
+We further fuse the integer subtraction into the Tensor Core mma (Matrix Multiply-Accumulate $D=AB+D$) instruction.
 We initialize the accumulator $D$ as the bias_int before MatMul computation of each quantization group, and will only need to subtract the bias_fp after the MatMul.
@@
 The above I2F simplification brings more than 20 TOPS performance improvement for the 512/4096/4096 (M/N/K) quantized MatMul computation on an A100 GPU.
-Figure 4: The GPU kernel software pipeline of group-wise W4A8/W8A8 quantized MatMul.
-It assumes perfect overlapping.
-G2S: load global to shared memory; S2R: load shared memory to register; MMA: matrix multiply-accumulation; I2F: integer to float conversion; deq: dequantize; acc: accumulate.
 End-to-end software pipeline of the quantized linear kernel on the GPU.
 Fig.3 shows the software pipeline of the quantized kernel.
-Beside the basic warp tile and block tile, we introduce the quantization group tile for the fast I2F and per-group scale multiplication.
+Besides the basic warp tile and block tile, we introduce the quantization group tile for the per-group quantized computation.
 It uses two output buffers for the output accumulation at register level, one for the per-group accumulation, and the other for the global accumulation.
@@
 As for the two-step dequantization, the first step is within the warp tile where each input element will subtract the zero-point before feeding into MMA, the second step is at the end of the group tile by multiplying the scale.
-We use the vectorized intrinsic to perform four int8 subtract in a single instruction (vsub4[27].
+We use the vectorized intrinsic to perform four int8 subtract in a single instruction (vsub4Lin et al.
+(2025).
 Besides, to improve the global memory loading efficiency, we prepack the memory layout of the weight tensor ahead-of-time to avoid the runtime permutation of the input elements.
-In general, this software pipeline can overlap the memory loading and computation, and the dequantization computation with SIMT Core and the MatMul computation with Tensor Core to the best, and minimizes the overhead of group-wise dequantization.
-Parallel execution of sub-problems of different bit-width.
+This software pipeline can overlap the memory loading, the dequantization computation with SIMT Core, and the MatMul computation with Tensor Core to the best, and minimizes the overhead of group-wise dequantization.
+3.4 Parallel Execution of Sub-problems of Different Bit-width
 As for the execution shown in Fig.1, MixLLM executes different sub-problems in parallel on the GPU with CUDA Graph.
@@
 We implement this function with the fused epilogue of MatMul to scatter the output to the corresponding indices, which is basically costless.
+3.5 Quantization-aware Roofline Discussion
+To systematically analyze the performance bounds of MixLLM quantization schemes, we extend the traditional Roofline model to account for the heterogeneous compute pipelines on NVIDIA GPUs.
+Using the NVIDIA A100 (80GB SXM) as our target architecture, we evaluate the system against its peak hardware limits: 2039 GB/s of High Bandwidth Memory (HBM2e) throughput, 624 TOPS of dense INT8 Tensor Core compute, and 19.5 TFLOPS of FP32 SIMT compute.
+Introducing sub-byte quantization and asymmetric zero-points fundamentally alters the arithmetic intensity (operations per byte) and introduces multiple competing compute ceilings across the execution units.
+Memory Traffic.
+From a memory bandwidth perspective, arithmetic intensity is governed by the quantized data footprint and associated metadata.
+For a group size of $G=128$, the W8A8 symmetric scheme fetches 8-bit weights, 8-bit activations, and a 16-bit scale, yielding an effective weight footprint of $1+2/G$ bytes per element.
+The W4A8 asymmetric scheme compresses weights to 4 bits, requiring a 16-bit scale and a 4-bit zero-point, yielding an effective footprint of $0.5+2.5/G$ bytes per weight.
+Consequently, W4A8 drastically shifts the memory-bound slope upward, allowing significantly higher attainable throughput in bandwidth-constrained regimes, such as small batch decoding.
+This motivates quantization with lower bits for weight in MixLLM.
+Compute Ceilings: Tensor Core vs.
+SIMT Dequantization.
+In the compute-bound regime, performance is dictated by the interplay between the Tensor Cores (handling the GEMM) and the SIMT cores (handling dequantization).
+For both W8A8 and W4A8, output scaling requires a transition from the Tensor Core’s int32 accumulators to the SIMT core’s float32 pipeline.
+For every group size of $G=128$, the Tensor Cores perform 128 INT8 multiply-accumulate (MAC) operations (256 operations total) to produce a single int32 partial sum.
+Before the FP32 scales ($s_{w}$ and $s_{a}$) can be applied, the SIMT cores must first issue an int32-to-float32 cast instruction.
+Therefore, the post-GEMM workload per group consists of this explicit cast instruction followed by the FP32 scale multiplications.
+Even with this added casting overhead, the application’s required operational ratio—256 INT8 operations per 3 SIMT instructions (cast + multiplies)—remains well below the A100 hardware’s provisioned ratio of 32:1 (624 INT8 TOPS vs.
+19.5 FP32 TFLOPS), thanks to the fast I2F and partial fusion methodology in MixLLM.
+Because the SIMT requirement is so light, this post-GEMM casting and scaling can be hidden behind the Tensor Core accumulation.
+This further motivates the group-wise 8-bit quantization than per-channel/per-token solutions.
+However, W4A8 introduces a heavier pre-GEMM dequantization step:
+$Y=((W_{q}-z)\times A_{q})\cdot s_{w}\cdot s_{a}$
+Before the Tensor Cores can execute the INT8 MACs, the 4-bit weights must be unpacked and the zero-point subtracted.
+By leveraging the vsub4 intrinsic, the hardware can perform 4 int8 subtractions within a single SIMT instruction, highly optimizing the $(W_{q}-z)$ computation.
+Despite this instruction-level parallelism, the SIMT integer pipeline must still continuously feed the Tensor Cores.
+Because this pre-GEMM workload scales at $O(N^{2})$ and competes for instruction fetch and register bandwidth concurrently with the $O(N^{3})$ Tensor Core math, it creates an auxiliary, lower compute ceiling.
+Thus, the effective peak compute for W4A8 is strictly lower than the absolute 624 TOPS achieved by W8A8.
+Intersection Point (Ridge Point) Analysis.
+The traditional Roofline ridge point defines the intersection between memory bandwidth and peak compute, marking the transition from a memory-bound to a compute-bound workload.
+For pure INT8 Tensor Core execution on the A100, this intersection occurs at roughly 306 OPs/Byte (624 TOPS / 2039 GB/s).
+Because the W4A8 scheme is constrained by the SIMT vsub4 dequantization ceiling rather than the peak Tensor Core ceiling, its horizontal Roofline ceiling drops.
+Consequently, the intersection between the W4A8 memory traffic line and its compute ceiling shifts to the left.
+This leftward shift indicates that W4A8 workloads saturate their maximum achievable compute at a lower arithmetic intensity than W8A8, entering the compute-bound regime earlier due to SIMT instruction overhead.
 4 Evaluation
-4.1 Setup
-As for MixLLM evaluation in this paper, we use 0%, 10%, 20%, 50% and 100% percent of 8-bit based on the 4-bit quantization, respectively.
-Meanwhile, we use 8-bit for activation quantization.
-Both the weight and activation are group-wise quantized with group size 128.
-The 4-bit part is asymmetric quantized and the 8-bit part (including that in weight) is symmetric, which is a good trade-off between accuracy and system efficiency.
-Note that any other bit-width percentage configuration can be used for real scenarios to trade-off memory usage, system efficiency and accuracy in practice.
-We enable GPTQ in MixLLM for the all models.
-We also apply clip search for all the models except for the 32B, 70B and 72B models as the clip search is time consuming for large models.
-Baselines and configurations.
-We compare MixLLM with the state-of-the-art (SOTA) quantization solutions of both weight-only and weight-activation methods.
-As for the weight only quantization, we compare MixLLM with the basic round-to-nearst (RTN) 4-bit and 5-bit quantization, and the production-level SOTA GPTQ [14] and AWQ[26].
-As for the weight-activation quantization, we compare MixLLM with the most widely used SmoothQuant [45] and the recent SOTA QoQ [27] and QuaRot[3] (of both W4A4 and W4A8).
-The 8-bit tensors are all symmetric quantized in all baselines and MixLLM.
-We also compare the perplexity with SqueezeLLM[21], OminiQuant[37], AffineQuant[29], Atom[50] and SpinQuant[28] according to their reported numbers.
-We make use of AutoGPTQ lib [5] (v0.7.1) to evaluate GPTQ, AutoAWQ lib [4] (v0.2.7) to evaluate AWQ, lmquant lib [33] (commit 9d62c5c) to evaluate SmoothQuant and QoQ, and the official repo to evaluat QuaRot.
-We enable the reorder trick for GPTQ evaluation, and use asymmetric and group size 128 for both GPTQ and AWQ.
-We follow the official configurations in lmquant to use 0.85/0.15 as the alpha/beta parameter for SmoothQuant, and 0.3/0.7 for QoQ.
-We use symmetric and per-channel/token quantization in QuaRot, following the configuration in its paper.
-We disable the KV quantization of QoQ and QuaRot in our experiments to make the comparison fair.
-Models and Datasets.
-We evaluate MixLLM and the baselines on a variety of widely used LLMs of different sizes, ranging from 0.5B to 72B.
-The models include Llama 3.1 8B and 70B [32], Llama 3.2 1B, Qwen2.5 0.5B, 1.5B, 7B and 32B [17], and Mistral 7B v0.3 [20].
-We use wikitext2 dataset [31] as the calibration set for GPTQ and MixLLM.
-We use the default pile dataset [34] as the calibration dataset for AWQ, SmoothQuant and QoQ, to enable their better performance.
-GPTQ, AWQ and MixLLM uses 128 samples with sequence length of 2048 for calibration.
-SmoothQuant and QoQ uses 64 samples with sequence length of 1024 for calibration (larger dataset results in OOM in our experiment).
-Metrics.
-As for the algorithm accuracy, we compare the perplexity (ppl) between all the baselines on wikitext2 and C4 [35] dataset.
-Meanwhile, we compare a set of popular downstream tasks on Llama 3.1 8B, Qwen2.5 7B, and Mixtral 7B v0.3, including BBH [39], GPQA [36], MMLU-Pro [41], MuSR [38], ARC challenge [6], and HellaSwag [49].
-We use lm-eval [15] for the downstream tasks evaluation, for which the task names are leaderboard_bbh, leaderboard_gpqa, leaderboard_mmlu_pro, leaderboard_musr, arc_challenge, and hellaswag.
-We use the average number for each of the tasks.
-We conduct the system experiments on NVIDIA A100 (80G) GPUs with CUDA 12.1. We use PyTorch 2.4.1 and transformers 4.45.2.
-4.2 Perplexity Evaluation
-Table 1: Perplexity evaluation ($\downarrow$) on wikitext2/c4 (gray for c4), sequence length 2048. NA means no support. Abn means the value is too large ($>10^{5}$). For MixLLM, pn means $n\%$ 8-bit.
+Table 1: Perplexity evaluation ($\downarrow$) on wikitext2/c4 (gray for c4), sequence length 2048. NA means no support. Abn means the value is too large ($10^{5}$). For MixLLM, pn means $n\%$ 8-bit.
 Llama 3.1/3.2 Qwen2.5 Mistral
@@
 W4A16 11.72/15.56 6.82/9.72 3.55/7.43 15.54/20.55 10.35/14.35 7.23/10.88 5.27/9.14 5.51/8.04
-RTN W5A16 10.15/13.25 6.40/9.15 3.16/9.52 13.61/18.17 9.52/13.38 6.95/10.53 5.09/8.99 5.38/7.91
-GPTQ W4A16 10.38/14.15 6.52/9.55 Abn/Abn 14.01/19.04 9.64/13.75 7.09/10.75 5.20/9.08 5.49/8.19
-AWQ W4A16 10.81/14.12 6.65/9.48 3.28/6.96 15.04/19.75 9.95/13.85 7.10/10.71 5.23/9.08 5.44/7.98
+Basic RTN W5A16 10.15/13.25 6.40/9.15 3.16/9.52 13.61/18.17 9.52/13.38 6.95/10.53 5.09/8.99 5.38/7.91
 SmoothQuant W8A8 9.89/12.91 6.34/9.08 2.92/6.77 13.84/18.40 9.63/13.49 7.17/10.85 5.12/9.04 5.35/7.88
-QoQ W4A8 Abn/Abn 6.64/9.49 3.49/7.07 Abn/Abn Abn/Abn 7.39/11.06 5.55/9.31 5.44/7.98
 W4A4 Abn/Abn 8.34/11.95 6.16/9.91 NA/NA Abn/Abn 8.15/12.05 6.26/9.98 5.83/8.50
 QuaRot W4A8 Abn/Abn 6.60/9.67 3.43/7.10 NA/NA Abn/Abn 7.03/10.68 5.23/9.10 5.40/7.99
+QServe W4A8 Abn/Abn 6.64/9.49 3.49/7.07 Abn/Abn Abn/Abn 7.39/11.06 5.55/9.31 5.44/7.98
 W4A8 (p0) 10.36/14.09 6.54/9.62 3.30/7.24 14.43/19.61 9.66/13.79 7.03/10.75 5.21/9.08 5.42/8.02
@@
 MixLLM W8A8 (p100) 9.76/12.75 6.25/8.97 2.81/6.68 13.12/17.60 9.28/13.14 6.86/10.45 5.02/8.96 5.32/7.84
-Tab.1 shows the perplexity on Wikitext2 and C4 dataset for the commonly used open source LLMs, of different baselines. It shows that:
-  • Using 4.4 bits of weights with MixLLM can achieve the similar accuracy with 5 bits RTN weight-only quantization, even with 8-bit activation quantization enabled in MixLLM.
+4.1 Setup
+As for MixLLM evaluation, we use 0%, 10%, 20%, 50% and 100% of 8-bit based on the 4-bit quantization, respectively.
+Meanwhile, we use 8-bit for activation quantization.
+Both the weight and activation are group-wise quantized with group size 128.
+The 4-bit part is asymmetric quantized and the 8-bit part (including that in weight) is symmetric.
+Similar to the recent work Lin et al.
+(2025); Liu et al.
+(2024), we enable GPTQ and clipping in MixLLM.
+Baselines and configurations.
+In Sec.4.2 and Sec.4.4, we compare to the pure PTQ methods of weight-activation quantization, neither the QAT nor the methods training the quantization parameters (e.g., training the rotation matrix like in SpinQuantLiu et al.
+(2024)).
+Specifically, we compare MixLLM with the SOTA PTQ weight-activation methods, including the most widely used SmoothQuant Xiao et al.
+(2023) and the recent SOTA QuaRotAshkboos et al.
+(2024) (of both W4A4 and W4A8) and QServe Lin et al.
+(2025).
+The 8-bit tensors are all symmetric quantized in all baselines.
+We use symmetric and per-channel/token quantization in QuaRot, following the setting in its paper.
+We follow the official configurations to use 0.85/0.15 as the alpha/beta parameter for SmoothQuant, and 0.3/0.7 for QServe.
+We disable the KV quantization of QuaRot and QServe in our experiments to make the comparison fair.
+In Appendix.4.6, We also compare the ppl with GPTQ Frantar et al.
+(2022), AWQLin et al.
+(2024), SqueezeLLMKim et al.
+(2024), OmniQuantShao et al.
+(2024), AffineQuantMa et al.
+(2024), AtomZhao et al.
+(2024) and SpinQuantLiu et al.
+(2024).
+Models and Datasets.
+We evaluate MixLLM and the baselines on Llama 3.1 8B and 70B Meta (Cited 2024), Llama 3.2 1B, Qwen2.5 0.5B, 1.5B, 7B and 32B Group (Cited 2024), and Mistral 7B v0.3 Jiang et al.
+(2023).
+We use wikitext2 Merity et al.
+(2017) as the calibration set for MixLLM, and the default pile dataset MIT-Han-Lab (Cited 2024) for SmoothQuant and QServe.
+MixLLM uses 128 samples with sequence length of 2048.
+SmoothQuant and QServe uses 64 samples with sequence length of 1024 to prevent OOM.
+Metrics.
+We compare the perplexity (ppl) between all the baselines on wikitext2 and C4 Raffel et al.
+(2020) datasets.
+We also compare a set of popular downstream tasks on Llama 3.1 8B, Qwen2.5 7B, and Mixtral 7B v0.3 through lm-eval Gao et al.
+(2024), including BBH Suzgun et al.
+(2023), GPQA Rein et al.
+(2023), MMLU-Pro Wang et al.
+(2024), MuSR Sprague et al.
+(2024), ARC challenge Boratko et al.
+(2018), and HellaSwag Zellers et al.
+(2019).
+We conduct the system experiments on NVIDIA A100 (80G) GPUs with CUDA 12.1, except for that Sec.4.3.2 discusses the performance on NVIDIA H100 GPU.
+We use PyTorch 2.4.1 and transformers 4.45.2.
+4.2 Perplexity Evaluation
+Tab.1 shows the perplexity on Wikitext2 and C4 dataset for the commonly used open source LLMs, of different baselines.
+It shows that: 1) Using 4.4 bits of weights with MixLLM can achieve the similar accuracy with 5 bits RTN weight-only quantization, even with 8-bit activation quantization enabled in MixLLM.
 This is mainly because MixLLM assigns the high-salience output channels with larger bit-widths than the uniform 5-bit solution.
-  • As for the weight-only quantization baselines, MixLLM W4.4A8 outperforms the production SOTA solutions GPTQ and AWQ, with only 10% more bit-width, and even with 8-bit activation quantization enabled in MixLLM.
-Meanwhile, the RTN W5A16 method also outperforms GPTQ and AWQ, which means a slightly larger bit-width can defeat the well tuned smaller bit-width easily.
-MixLLM W4.4A8 benefits from the larger bits on the top 10% output features with high salience.
-  • As for the weight-activation quantization baselines, MixLLM W4.4A8 shows a comparable accuracy with SmoothQuant with much smaller bit-width (60% of that in SmoothQuant).
-MixLLM W4.4A8 shows better accuracy than QoQ and QuaRot with only 10% larger bit-width.
+2) As for the weight-activation quantization baselines, MixLLM W4.4A8 shows a comparable accuracy with SmoothQuant with much smaller bit-width (60% of that in SmoothQuant).
+MixLLM W4.4A8 shows better accuracy than QuaRot and QServe with only 10% larger bit-width.
 It shows MixLLM achieves a good balance of memory consumption and accuracy.
-  • Note that MixLLM W8A8 quantization shows nearly lossless performance compared to the float16 baseline. This is part of the motivation that MixLLM uses group-wise quantization for the activation.
-Table 2: PPL (wikitext2) comparison with the reported numbers in the related works.
-LLaMA 2 FP16
-SqueezeLLM
-W4A16 0.45%
-OminiQuant
-W4A16/W4A4
-AfineQuant
-W4A16/W4A4
-Atom
-W4A4 128 outliers
-SpinQuant
-W4A8
-MixLLM
-W4.4A8
-7B 5.47 5.57 5.58/14.26 5.58/12.69 6.03 5.7 5.55
-13B 4.88 4.96 4.95/12.30 4.95/11.45 5.27 5.0 4.93
-Comparison with More Related Works We compare MixLLM with more recent quantization works according to the reported numbers in their papers (Tab.2), showing that MixLLM achieves superior accuracy to a broad range of related works with similar memory consumption.
+3) Note that MixLLM W8A8 quantization shows nearly lossless performance compared to the float16 baseline.
+Besides, the MixLLM W4A8 also outperforms the SOTA QuaRot and QServe for many cases, due to using group-wise quantization for the activation in MixLLM rather than the per-token method in QuaRot and QServe.
+This is part of the motivation that MixLLM uses group-wise quantization for the activation.
+Figure 5: The speedup of two types of single linear layers over torch FP16 baseline on A100 GPU.
 4.3 System Performance
-We have evaluated MixLLM for the single linear layer of token number ranging from 1 to 1024 with in features 4096 and out features 4096/14336, and compared it with the SOTA W4A16 (TRT-LLM) and QoQ [27], shown in Fig.5.
-It also shows MixLLM kernel performance of different percent of 8-bits (W4A8 0% 8-bit, W4.4A8 10% 8-bit, and W8A8 100% 8-bit).
-It shows that:
-  • MixLLM outperforms the float16 counterpart for all token numbers, achieving 1.90$\times$, 2.75$\times$, and 1.88$\times$ averaged speedup with MixLLM W4A8, W8A8, and W4.8A8 respectively.
-  • MixLLM outperforms the SOTA W4A16 solution, achieving 1.26$\times$, 1.78$\times$, and 1.25$\times$ averaged speedup with MixLLM W4A8, W8A8, and W4.8A8 respectively.
-  • MixLLM achieves similar performance with QoQ with similar bit-width, achieving 0.99$\times$, 1.39$\times$, and 0.99$\times$ averaged speedup with MixLLM W4A8, W8A8, and W4.8A8 respectively.
-Note that MixLLM has better accuracy than QoQ (Tab.1, Tab.3).
-Figure 5: The speedup of two types of single linear layers over torch float16 baseline on the A100 GPU.
-4.4 Downstream Tasks Evaluation
+4.3.1 Performance on A100 GPU
+[…]
+Figure 6: Speedup of MixLLM W8A8 of Qwen 2.5 7B over float16 baseline on A100 GPU.
+[…]
+4.3.2 Performance on H100 GPU
+Figure 7: The speedup of two types of single linear layers over W4A16 baseline on H100 GPU.
+We evaluate the performance of single linear layers on the NVIDIA H100 GPU, comparing the MixLLM W4A8 kernel against the W4A16 baseline (Fig.
+7).
+To the best of our knowledge, an H100-optimized implementation of the QoQ solution Lin et al.
+(2025) is currently unavailable.
+The W4A16 baseline was established by identifying the optimal configuration through exhaustive tuning of the official CUTLASS group-wise and asymmetric-quantized W4A16 examples.
+Our analysis considers both a naive 4-bit layout and an advanced shuffled 4-bit layout.
+Results indicate that MixLLM consistently outperforms the W4A16 solutions across all configurations.
+Specifically, for an output feature size of 4096, MixLLM achieves average speedups of $1.81\times$ and $1.39\times$ over the naive and shuffled layouts, respectively; for a size of 14336, these speedups are $1.91\times$ and $1.34\times$.
 Table 2: Downstream tasks evaluation ($\uparrow$) on Llama-3.1-8B/Qwen2.5-7B/Mistral-7B-v0.3. The above is the average of the three models. BBH is 3 shot, MMLU pro is 5 shot, and others are zero shot.
@@
 78.92/78.94/80.43
-GPTQ W4A16
-47.08
-46.28/51.97/43.00
-30.98
-30.81/34.59/27.53
-33.66
-31.20/41.02/28.75
-41.83
-39.06/44.64/41.78
-51.37
-53.24/48.98/51.88
-78.46
-77.85/78.10/79.44
-AWQ W4A16
-47.62
-45.59/52.71/44.57
-30.86
-29.76/33.17/29.65
-34.31
-31.57/42.78/28.59
-40.36
-38.27/42.37/40.44
-51.17
-52.39/50.60/50.51
-78.77
-78.29/78.35/79.68
 SmoothQuant W8A8
@@
 78.88/78.48/80.24
-QoQ W4A8
-45.78
-40.98/51.23/45.14
-30.02
-28.99/32.50/28.56
-32.84
-28.16/41.72/28.63
-39.92
-37.60/41.59/40.57
-50.54
-51.28/49.15/51.19
-78.10
-76.90/77.52/79.89
 QuaRot W4A4
@@
 77.83/77.84/79.98
+QServe W4A8
+45.78
+40.98/51.23/45.14
+30.02
+28.99/32.50/28.56
+32.84
+28.16/41.72/28.63
+39.92
+37.60/41.59/40.57
+50.54
+51.28/49.15/51.19
+78.10
+76.90/77.52/79.89
 MixLLM W4A8
@@
 78.98/78.88/80.40
-Tab.3 shows the accuracy of the downstream tasks on three popular LLMs. The result shows that:
-  • MixLLM W4.4A8 outperforms all the 4-bit weight quantizations, with only 10% more weight consumption.
-For example, for the MMLU-Pro task, the average metric of MixLLM W4.4A8 is improved by 1.69, 6.93, and 0.93 over QoQ, QuaRot W4A4, and QuaRot W4A8, respectively.
-  • MixLLM W8A8 is nearly lossless, showing higher accuracy than SmoothQuant. This comes from the group-wise quantized activation of MixLLM.
-4.5 Detailed Analysis
-4.5.1 Ablation Study
+4.4 Downstream Tasks Evaluation
+We first evaluate GSM8K Cobbe et al.
+(2021) on Qwen2.5 7B model, to validate the quantization efficacy on long-reasoning tasks.
+Using MixLLM W4.4A8 quantization, the strict-match metric only drops from 0.8 (the float16 model) to 0.792, only a 0.008 drop.
+Besides this, we evaluate a large number of the downstream tasks on three popular LLMs, shown in Tab.2.
+The result shows that: 1) MixLLM W4.4A8 outperforms all the 4-bit weight quantizations, with only 10% more bit-width.
+For example, for the MMLU-Pro task, the average metric of MixLLM W4.4A8 is improved by 1.69, 6.93, and 0.93 over QServe, QuaRot W4A4, and QuaRot W4A8, respectively.
+2) MixLLM W8A8 is nearly lossless, showing higher accuracy than SmoothQuant.
+This comes from the group-wise quantized activation of MixLLM.
+4.5 Ablation Study
 Figure 8: The perplexity (wikitext2) of Llama 3.1 8B model with different configurations.
@@
 Finally, applying GPTQ and clipping can further improve the accuracy.
-4.5.2 High Precision Distribution
-Fig.2 shows the percentage of 8-bit out features in each of the linear layers of Llama 3.1 8B, with 10% global 8-bit out features searched by MixLLM (i.e., W4.4A8).
-It shows that high-salient (i.e., 8-bit) features are distributed very differently in different linear layers.
-Specifically, the v_proj and down_proj layers show much higher percentage of high-salient features than other layers, for which Tab.4 shows the average percentage of different classes of linear layers.
-Table 4: The average percentage of 8-bit out features in the seven classes of linear layers in Llama 3.1 8B, with 10% global 8-bit out features in MixLLM.
-layer q_proj k_proj v_proj o_proj gate_proj up_proj down_proj
-avg 8-bit percent (%) 3.93 12.36 71.22 18.70 0.73 1.46 53.82
-4.5.3 One-pass vs. Progressive Search
-As described in Sec.3.2, MixLLM searches the high-salience features within a single pass.
-We have tried the progressive procedure on Llama 3.1 8B and Mistral 7B models, which identifies smaller portions of the high-salience features iteratively.
-Results show that the accuracy is the same to the one-pass method to two decimal places.
-However, the progressive method shows much higher search time due to the repeated procedure.
-The one-pass method takes 7 minutes for each of the two models to search 10% high-salience features, while the progressive method that searches 2% high-salience iteratively takes 30 minutes to find top 10% features.
-4.5.4 Overhead of Global Precision Search
-Table 5: The overhead of global precision search in MixLLM.
+Table 3: PPL (wikitext2) comparison with the reported numbers in the related works.
+Model FP16
+GPTQ
+W4A16
+AWQ
+W4A16
+SqueezeLLM
+W4A16 0.45%
+OmniQuant
+W4A16/W4A4
+AfineQuant
+W4A16/W4A4
+Atom
+W4A4 128 outliers
+SpinQuant
+W4A8
+MixLLM
+W4.4A8
+LLaMA 2 7B 5.47 5.59 5.60 5.57 5.58/14.26 5.58/12.69 6.03 5.7 5.55
+LLaMA 3 8B 6.14 6.46 6.55 - - - 7.57 6.5 6.32
+4.6 Comparison with More Related Work
+Table 4: PPL of Atom and MixLLM with similar bit-width.
+Models Llama 2 7B Llama 2 13B
+Atom W4.4A8 5.64 5.03
+MixLLM W4.4A8 5.54 4.93
+Comparison with Atom.
+Tab.4 shows the perplexity of Atom and MixLLM with the similar bit-width (i.e., W4.4A8).
+We use 512 outliers in Atom so its weight is around 4.4-bit on average (using Atom’s opensourced code with git commit 7e3618b).
+It shows that MixLLM has much better accuracy than Atom.
+We also evaluated the kernel performance of Atom, showing that our W4.4A8 kernel achieves 1.56x speedup than Atom’s W4A8 kernel for the linear layer with sequence length 1024.
+This means our proposed memory and computation pipeline has much better performance than the related mixed-precision work.
+Comparison with SliM-LLM.
+Tab.5 shows the perplexity of SliM-LLM W4A16 (it only supports weight-only) and MixLLM with different bit-width.
+SliM-LLM W4A16 is evaluated using the opensourced code.
+It is the mix of 3/4/5-bits in its code, and the precision is searched within each single layer, rather than using the global precision search.
+Note that MixLLM can support the mix of any bit-width.
+We evaluated MixLLM with the mix of 4-bit/6-bit (W4.2A8, 4.2-bit on average with 90% 4-bit and 10% 6-bit), and the mix of 3-bit/4-bit/6-bit (W3.9A8, 3.9-bit on average with 30% 3-bit, 60% 4-bit, and 10% 6-bit) in this table.
+We notice even MixLLM W4A8 (uniform 4-bit for the weight) can defeat SliM-LLM W4A16.
+This is understandable because SliM-LLM itself focuses on 2-bit and 3-bit optimization, and its paper’s main body does not show any 4-bit results.
+In contrast, MixLLM W3.9A8 can defeat MixLLM W4A8.
+This demonstrates that MixLLM’s global precision search can better allocate bit-width to the important weight elements to achieve the nearly lossless results.
+Note that the percentage of bit-widths is determined intuitively here, so that it may not be the optimal percentage to achieve the optimal accuracy.
+How to determine the portion of different bit-width can be a new research problem.
+Table 5: PPL (wikitext2) of SliM-LLM and MixLLM.
+Models Llama Qwen2.5
+3.2 1B 3.1 8B 0.5B 1.5B 7B
+SliM-LLM W4A16 10.80 6.55 14.85 9.68 7.02
+MixLLM W3.9A8 10.20 6.50 13.68 9.51 6.97
+MixLLM W4A8 (uniform) 10.36 6.54 14.43 9.66 7.03
+MixLLM W4.2A8 10.07 6.43 13.53 9.45 6.93
+MixLLM W4.4A8 10.05 6.42 13.42 9.44 6.92
+We compare MixLLM with more recent quantization works according to the reported numbers in their papers (Tab.3), showing that MixLLM achieves superior accuracy to a broad range of related works with similar memory consumption.
+4.7 Overhead of Global Precision Search
+Table 6: The overhead of global precision search.
 Models Llama 3.1 Mistral Qwen2.5
@@
 Considering that the quantization only needs to be performed once, the searching algorithm is practical for the real workloads.
+In contrast, SliM-LLM takes more than 3 hours to determine the precision of Llama 3.1 8B model and more than 4 hours for Qwen2.5 7B model according to our experiments.
+4.8 High Precision Distribution
+Fig.2 shows the percentage of 8-bit out features in each of the linear layers of Llama 3.1 8B, with 10% global 8-bit out features searched by MixLLM (i.e., W4.4A8).
+It shows that high-salient (i.e., 8-bit) features are distributed very differently in different linear layers.
+Specifically, the v_proj and down_proj layers show much higher percentage of high-salient features than other layers, for which Tab.7 shows the average percentage of different classes of linear layers.
+Table 7: The average percentage of 8-bit out features in the seven classes of linear layers in Llama 3.1 8B, with 10% global 8-bit out features in MixLLM.
+layer (xx_proj) q k v o gate up down
+avg 8-bit (%) 3.93 12.36 71.22 18.70 0.73 1.46 53.82
+4.9 One-pass vs. Progressive Search
+As described in Sec.3.2, MixLLM searches the high-salience features within a single pass rather than iteratively identifying the high-salience parts in a smaller step, as we observe the single-pass method show similar results with the iterative method and saves a lot of computation overhead than the latter.
+We have tried the progressive procedure on Llama 3.1 8B and Mistral 7B models, which identifies smaller portions of the high-salience features iteratively.
+Results show that the accuracy is the same to the one-pass method to two decimal places.
+However, the progressive method shows much higher search time due to the repeated procedure.
+The one-pass method takes 7 minutes for each of the two models to search 10% high-salience features, while the progressive method that searches 2% high-salience iteratively takes 30 minutes to find top 10% features.
+4.10 Working with KV Quantization
+Table 8: PPL of enabling KV Quantization in MixLLM.
+Models Llama 3.1 8B Qwen2.5 7B
+MixLLM W4.4A8 w/o KV quant 6.42 6.92
+MixLLM W4.4A8 KV8 6.42 6.96
+It is straightforward to apply any KV quantization technology together with MixLLM’s weight-activation quantization, as KV quantization is orthogonal to the weight-activation quantization.
+Tab.8 shows the perplexity (wikitext2) of Llama 3.1 8B and Qwen 2.5 7B for enabling and disabling KV quantization.
+It shows that the 8-bit KV quantization is nearly lossless upon MixLLM.
+4.11 Calibration Dataset for Salience Search
+We use 128 samples for the calibration to search the high-salience channels in MixLLM, which is a common configuration in the existing solutions.
+We also find that a smaller dataset can still identify the high-salience channels in MixLLM.
+The perplexity of using different samples for MixLLM salience search are shown in Tab.9.
+This shows that shrinking the sample size from 128 to 16 for the global salience search has negligible effect to the accuracy.
+Table 9: PPL (wikitext2) when using different number of samples for salience search (W4.4A8 quantized).
+#samples 128 64 32 16
+Llama 3.1 8B 6.42 6.42 6.42 6.42
+Qwen2.5 7B 6.92 6.92 6.92 6.93
+We also use different dataset families between salience searching and perplexity evaluation to evaluate the effect of input distribution shift.
+When using c4 dataset for salience search and evaluate the perplexity on wikitext2 (W4.4A8 quantized), the Llama 3.1 8B and Qwen2.5 7B has the same perplexity with that using wikitext2 for the calibration.
+This shows that changing the calibration dataset does not affect the accuracy for the same task.
+In another word, when the data distribution between the calibration dataset and real input is different, MixLLM’s salience search algorithm still works well.
 5 Summary
 We have presented MixLLM, achieving high accuracy with low memory consumption and high system efficiency with the rarely explored optimization space of mixed-precision quantization between output features.
-MixLLM identifies the salience of each output feature according to the loss distance estimation w.r.t.
-the global model loss rather than local layer loss.
-By assigning larger bit-width to the features need it most, MixLLM achieves the superior accuracy to SOTA with low memory consumption.
+MixLLM identifies the salience of each output feature according to the loss w.r.t.
+the global model rather than each single layer.
+By assigning larger bit-width to the features need it the most, MixLLM achieves the superior accuracy to SOTA with low memory consumption.
 The sub-problems of different bit-widths are disjoint and can run in parallel efficiently on the GPU.
-We have identified the sweet spot of the quantizaiton configuration that is friendly to both accuracy and system efficiency.
-To address the challenge of system efficiency, we design the two-step dequantization to enable using int8 Tensor Core computation and the fast integer-float conversion to reduce the dequantization overhead.
-We have designed the end-to-end software pipeline to overlap the memory access, the dequantization computation with SIMT Core and the MatMul with Tensor Core.
+We have identified the sweet spot of the quantization configuration that is friendly to both accuracy and system efficiency.
+To address the challenge of system efficiency, we design the two-step dequantization to enable using int8 Tensor Core and the fast integer-float conversion to reduce the dequantization overhead.
+We have designed the end-to-end software pipeline to overlap the memory access, the dequantization computation with SIMT Core and the MatMul with Tensor Core to the best.
 Experiment results show that MixLLM achieves superior accuracy to existing works and state-of-the-art system efficiency with low memory cost.